### v1

# Descrip3D: Enhancing Large Language Model-based 3D Scene Understanding with Object-Level Text Descriptions

Abstract
Understanding 3D scenes goes beyond simply recognizing objects; it requires reasoning about the spatial and semantic relationships between them. Current 3D scene-language models often struggle with this relational understanding, particularly when visual embeddings alone do not adequately convey the roles and interactions of objects. In this paper, we introduce Descrip3D, a novel and powerful framework that explicitly encodes the relationships between objects using natural language. Unlike previous methods that rely only on 2D and 3D embeddings, Descrip3D enhances each object with a textual description that captures both its intrinsic attributes and contextual relationships. These relational cues are incorporated into the model through a dual-level integration: embedding fusion and prompt-level injection. This allows for unified reasoning across various tasks such as grounding, captioning, and question answering, all without the need for task-specific heads or additional supervision. When evaluated on five benchmark datasets, including ScanRefer, Multi3DRefer, ScanQA, SQA3D, and Scan2Cap, Descrip3D consistently outperforms strong baseline models, demonstrating the effectiveness of language-guided relational representation for understanding complex indoor scenes.
1 Introduction
Recent advances in large language models (LLMs) have significantly transformed human-computer interaction by equipping machines with unprecedented capabilities in language understanding, reasoning, and open-ended dialogue. These models have demonstrated remarkable success in a wide range of domains, including information retrieval, code generation, and multi-modal learning [17, 35, 60, 28, 49, 42]. Building on this progress, researchers have begun to extend LLMs to 3D scene understanding [33, 57, 37], aiming to empower models with the ability to interpret and reason about complex visual and spatial contexts within real-world environments.
In particular, indoor scenes present unique challenges for 3D scene-language modeling, including dense layouts, ambiguous viewpoints, and complex spatial relationships. Earlier methods [4, 11] were typically designed for specific tasks such as grounding or captioning. In contrast, recent efforts [20, 22] explore 3D multimodal large language models (MLLMs) that integrate language with point clouds and multiview images to jointly support a wide range of tasks, such as 3D visual grounding, dense captioning, and question answering, using a single trained model. Models like Chat-Scene [22] adopt object-centric pipelines by associating 2D/3D embeddings with object identifiers and feeding them into the language model. However, these approaches still show limited performance on tasks requiring spatial or semantic reasoning across multiple objects, primarily because they lack explicit modeling of inter-object relationships and underutilize the language understanding capabilities of LLMs. Although recent graph-based methods [56] attempt to address this, they introduce additional training complexity and still fall short in performance.
Figure 1: An example of injecting object-level text descriptions during conversation. Providing these descriptions significantly improves the model’s accuracy and reasoning performance.
As illustrated in Figure 1, we address this limitation by introducing a novel text-based relational modality into the 3D scene language pipeline. Specifically, we augment each object with a natural language description that captures both its intrinsic attributes (e.g., color, material) and its contextual relationships to nearby objects (e.g., ’next to the table’, ’under the chair’), generated by prompting a vision-language model with multi-view images. We use a dual-level integration strategy to incorporate the relational descriptions into our model. These descriptions are embedded using a lightweight text encoder and fused with the object’s visual embeddings to create a unified multi-modal representation. In addition to embedding-level fusion, the descriptions are also injected into the prompt to support relational understanding in downstream tasks.
This dual-level integration, at both the representation and the language interface levels, allows the model to leverage the prior knowledge of the LLM about the spatial and functional relations learned from the language. Crucially, this approach does not introduce architectural changes, does not introduce task-specific heads, and does not require additional human annotation. It is modular, generalizable, and easy to integrate into existing object-centric pipelines.
Empirically, our method yields consistent gains across five standard 3D scene language benchmarks. ScanRefer [4], Multi3DRefer [58], Scan2Cap [11], ScanQA [2], and SQA3D [34]. Improvements are especially pronounced in tasks involving complex grounding or multi-object reasoning, highlighting the benefit of explicit text-based relational cues in enhancing 3D scene understanding.
Our contributions are as follows.
  • We identify and address a fundamental limitation in existing 3D scene-language models: the absence of explicit modeling of inter-object relationships, which hinders multi-object spatial and semantic reasoning.
  • We introduce a novel object-centric textual modality that encodes both intrinsic attributes and contextual relationships of objects through automatically generated natural language descriptions.
  • We propose a dual-level integration framework that incorporates these descriptions both at the embedding fusion level and within language prompts, enabling seamless and architecture-agnostic enhancement of existing LLM-based pipelines.
  • We validate our approach across five standard 3D scene-language benchmarks, consistently surpassing both expert-designed and LLM-based baselines, especially in tasks demanding relational understanding and compositional reasoning.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 details our proposed approach. Section 4 presents the experimental results and analysis. Finally, Section 5 concludes the paper.
2 Related work
2.1 3D Vision-Language Understanding
3D vision-language understanding combines 3D spatial perception with natural language to interpret and interact with complex environments. Earlier work focused on single-task models for visual grounding [30], dense captioning [53], and question answering [29]. For example, models such as ScanRefer [4], ReferIt3D [1], and Multi3DRefer [58] focus on grounding by locating target objects based on descriptions of natural language. In terms of dense captioning, models such as the 3DVG-Transformer [59] and Vote2Cap-DETR++ [9] work to both localize and describe objects within a 3D environment. Furthermore, question-answer models like ScanQA [2] and SQA3D [34] extend the task space by incorporating a broader understanding of the scene to answer general or context-dependent questions.
Recent research is increasingly focusing on multi-task frameworks that unify these objectives to enhance representation sharing and robust performance. For example, methods such as 3DJCG [3] and D3Net [5] unify 3D visual grounding and dense captioning through shared encoders and task-specific heads, leveraging complementary supervision across tasks to improve performance. Likewise, 3D-VisTA [62] and 3D-VLP [27] employ large-scale pre-training to learn generalizable 3D vision-language representations, enhancing performance across diverse downstream tasks. However, their reliance on task-specific heads and visual embeddings treats language as auxiliary, lacking explicit modeling of semantic object relationships.
2.2 Multi-modal Large Language Models
LLMs have demonstrated impressive capabilities in reasoning, generalization, and multi-turn dialogue, sparking interest in extending their input space to include 3D data [33]. In the 3D vision language domain, models such as PointLLM [48], ImageBind-LLM [18], Point-Bind [16], Ulip [50], and CG3D [19] extend LLMs to handle 3D data by aligning the embeddings of the point cloud with the language embedding space. These methods enable object-level understanding through joint representations of geometric and linguistic information, facilitating tasks such as 3D object captioning and question answering. However, they primarily operate on isolated object input and lack mechanisms for encoding explicit relationships between objects, which are essential for capturing scene-level semantics.
Multimodal large language models (MLLMs) have recently been extended to 3D scene understanding to support open-ended tasks such as question answering. Approaches such as 3D-LLM [20] map the embeddings of point clouds into the language space using position embeddings, allowing basic multimodal reasoning but lacking fine-grained localization and interactive grounding. Chat-3D [44] enhances the 3D dialogue by using object-centric prompts and a three-stage training scheme to align the characteristics and relations of objects with language models. However, its architecture restricts the model’s attention to individually selected objects, limiting its ability to reason in broader contexts of the scene. Chat-Scene [22] further advanced this line by incorporating object identifiers and aligning them with 2D and 3D visual embeddings, enabling strong performance on scene-level tasks such as question answering and visual grounding. Yet, they still lack explicit mechanisms for modeling inter-object relationships, limiting relational and spatial reasoning.
2.3 Explicit Object Relationship Representation
Although prior models lack explicit mechanisms for encoding inter-object relationships, recent efforts have begun to address this limitation through graph-based or language-based modeling. ConceptGraph [15] constructs scene-level concept graphs by aligning 3D object embeddings with textual descriptions using vision-language models, enabling semantic relationship modeling and open-vocabulary generalization. HOV-SG [45] advances this direction by generating holistic object-scene graphs from 3D scans, capturing spatial and semantic relations across the scene. However, such graph-based representations often require long token sequences and suffer from high inference costs in large-scale scenes. To address this, 3DGraphLLM [56] incorporates knowledge graph embeddings into object representations, encoding pairwise relationships to support relational reasoning. Yet, this approach introduces significant combinatorial overhead by enumerating object-object triplets and suffers from error propagation in graph construction. Notably, the addition of explicit graph structures in 3DGraphLLM leads to a performance drop on downstream question answering tasks, suggesting that current graph-based formulations may hinder rather than help relational reasoning in practice. Scene-LLM [14] instead leverages LLMs with scene-level prompts to generate global summaries and captions, but lacks fine-grained object-level relational modeling. Thus, explicitly and efficiently representing object relationships in 3D scenes remains an open and challenging problem.
3 Method
3.1 Overview
Our method empowers LLMs to perform precise and context-aware reasoning over complex 3D scenes by augmenting object representations with rich, relational textual descriptions. We introduce a powerful text-based relational modality that explicitly encodes both spatial and semantic relationships between objects. Our framework integrates pre-trained 3D and 2D visual encoders, a text encoder, and an LLM, and represents each object as a fused token embedding that combines geometric structure, visual appearance, and relational context. These relational cues are injected into both the object-level embeddings and directly into the LLM prompt, enabling fine-grained multimodal understanding and interaction. This dual-level integration significantly improves the ability of LLM to reason about object relationships, supporting downstream tasks such as grounding, captioning, and question answering with high accuracy and interpretability. An overview of our pipeline is shown in Figure 2.
Figure 2: Overall model architecture. We propose a novel and powerful method that explicitly models inter-object relationships by integrating relational text descriptions into object-centric scene representations via a dual-level strategy. From a 3D scan, we extract object proposals and encode their geometry and appearance using pretrained 2D and 3D encoders. Each object is enriched with a natural language description capturing both intrinsic attributes and spatial relations to nearby objects. These descriptions guide scene understanding through: (1) embedding-level fusion with visual features to enhance object representations, and (2) prompt-level injection of queried object descriptions to enhance object-specific relational reasoning. The resulting multimodal tokens enable high-level reasoning for 3D grounding, dense captioning, and question answering. Our design equips the model with both localized and contextual spatial semantics, significantly improving relational reasoning.
3.2 Object-Centric Scene Representation
We decompose each scene into discrete object proposals using Mask3D [39] first, producing a set of segmented object point clouds $\{\mathbf{P}_{1},...,\mathbf{P}_{n}\}$. Each object $\mathbf{P}_{i}\in\mathbb{R}^{m_{i}\times 6}$ includes XYZ coordinates and RGB values. Here, $m_{i}$ refers to the number of points in the proposal of the $i$th object. Each resulting object is treated as a fundamental unit for understanding the scene. To support unambiguous reference and interaction, we adopt the design introduced in Chat-Scene [22] to associate each object with a unique learnable identifier token $<\!OBJ_{i}\!>$. These tokens are integrated into the tokenizer vocabulary and embedded alongside the embeddings of the object.
3.3 Visual Embedding Extraction
For each detected object, we extract geometric and visual embeddings from 3D point clouds and multi-view images. To capture 3D geometry, we use the pre-trained Uni3D [61] encoder, which provides high-quality spatial embeddings by processing each object’s segmented point cloud $\mathbf{P}_{i}$, producing an embedding vector $\mathbf{Z}^{v_{p}}_{i}\in\mathbb{R}^{1\times d}$. Uni3D is chosen for its strong performance in capturing fine-grained geometric structure across diverse 3D scenes. For 2D appearance, we employ the pre-trained DINOv2 [36] vision transformer to extract embeddings from multi-view images. The 3D mask of each object is projected onto these images, and the corresponding regions are cropped from the embedding maps and aggregated into a 2D visual embedding $\mathbf{Z}^{v_{f}}_{i}\in\mathbb{R}^{1\times d}$. DINOv2 is used for its ability to extract semantically rich embeddings from high-resolution visual input.
3.4 Text-Based Relational Modality
While the embeddings extracted from both 2D multi-view images and 3D point clouds are powerful, these embeddings are inherently localized to individual objects. Specifically, 3D embeddings capture geometric and spatial attributes of each object in isolation, while 2D embeddings, though derived from full images, are ultimately aggregated per object and offer only limited relational context, making their relational awareness implicit and limited.
As a result, the visual representation lacks explicit modeling of inter-object interactions, limiting its capacity for comprehensive scene understanding and high-level reasoning. To overcome this limitation, we introduce a novel textual modality that encodes relational information in natural language. These descriptions explicitly capture how each object is situated within the scene and how it interacts with others, injecting rich contextual knowledge that is often inaccessible to vision-based embeddings alone.
Our approach introduces this modality through a dual-level integration: (1) by encoding and fusing the text descriptions into the object representations and (2) by injecting them into the prompt space of the language model for enhanced reasoning. This design allows our model to take advantage of both low-level visual embeddings and high-level relational cues, ultimately improving performance on tasks that require contextual understanding.
Relational Description Generation.
We generate natural language descriptions for each object by utilizing a vision-language model and apply it to multi-view RGB images. For each image, we project 3D object masks onto the 2D plane. This allows us to identify key objects located in the central area of the view, as well as all visible objects in each image. To ensure that the model correctly grounds each object, we overlay the object names (e.g., "couch", "shelf") on the image at their projected centers. We then ask the vision-language model to describe the relationship between each key object and all other visible objects in the image. This results in descriptions at the object level such as: “There is a curtain in the room. The curtain is covering the window, and it is also close to a table.”
We generate one description per object identifier, skipping any object that has already been described in a previous view to enhance efficiency. In this way, each object receives a single textual description when it appears at the center of at least one multi-view image, resulting in a complete set of object-level relational descriptions for the scene.
Relational Descriptions Encoding.
To encode the generated relational descriptions into the model, we use a Sentence-Transformer model [38] to convert each object’s description into a fixed-dimensional embedding vector $\mathbf{Z}^{v_{t}}_{i}$. This representation is subsequently fused into the object representation, complementing the 2D visual embeddings $\mathbf{Z}^{v_{f}}_{i}$ and the 3D geometric embeddings $\mathbf{Z}^{v_{p}}_{i}$.
Relational Augmented Prompting
In addition to token-level fusion, we leverage a prompt-level strategy that delivers relational priors directly to the language model. When a query mentions a specific object by name or identifier, we prepend the description of that object to the input prompt. To avoid semantic ambiguity and ensure consistent grounding, we replace all object names within the descriptions with their corresponding object identifiers. These object identifiers are learned alongside the language model during training, allowing the LLM to associate them with specific objects in the scene. This strategy eliminates confusion caused by duplicate object names (for example, multiple ’chairs’) and ensures that relational signals remain precisely anchored to the intended objects. The prompt-level text injection allows the LLM to access relevant contextual cues before reading the user’s question, improving reasoning performance.
3.5 Token Construction and Training Strategy
To enable unified reasoning across geometry, appearance, and language, all three modalities are projected into a shared token space compatible with the language model. Specifically, we employ separate linear projection heads $f_{p}$, $f_{v}$, and $f_{t}$ to map the original embeddings $\mathbf{Z}^{v_{p}}_{i}$, $\mathbf{Z}^{v_{f}}_{i}$, and $\mathbf{Z}^{v_{t}}_{i}$ into token embeddings $\mathbf{F}^{v_{p}}_{i}$, $\mathbf{F}^{v_{f}}_{i}$, and $\mathbf{F}^{v_{t}}_{i}$:
$\mathbf{F}^{v_{p}}_{i}$ $=f_{p}(\mathbf{Z}^{v_{p}}_{i}),$ $\mathbf{F}^{v_{f}}_{i}$ $=f_{v}(\mathbf{Z}^{v_{f}}_{i}),$ $\mathbf{F}^{v_{t}}_{i}$ $=f_{t}(\mathbf{Z}^{v_{t}}_{i}).$ (1)
Each object is represented by a fused token embedding constructed by concatenating its learned identifier token, its 3D geometry embeddings, 2D appearance embeddings, and textual description embeddings as
$\mathbf{F}_{i}=\texttt{Concat}(\mathbf{OBJ_{i}},\mathbf{F}^{v_{p}}_{i},\mathbf{F}^{v_{f}}_{i},\mathbf{F}^{v_{t}}_{i}).$ (2)
[…]
The model is optimized end-to-end using the cross-entropy loss over response tokens. The training objective is formulated as
$\mathcal{L}(\theta)=-\sum_{i=1}^{k}\log P\left(s_{i}^{\text{res}}\mid s_{[1,\dots,i-1]}^{\text{res}},\,s^{\text{prefix}}\right),$ (3)
where $k$ is the length of the response, $s_{i}^{\text{res}}$ is the generated target response, and $s_{[1,\dots,i-1]}^{\text{res}}$ is the previous $i-1$ tokens in the response. $\theta$ denotes the trainable parameters.
4 Experiments
4.1 Datasets and Metrics
Datasets.
To evaluate the generalizability and task effectiveness of our approach, we conduct experiments on five benchmark datasets spanning four major 3D vision-language tasks. These include: ScanRefer [4] for single-object localization via visual grounding, Multi3DRefer [58] for compositional grounding of multiple targets, Scan2Cap [11] for 3D-aware caption generation, and both ScanQA [2] and SQA3D [34] for contextual question answering within 3D scenes. All five datasets are derived from the ScanNet dataset [13], a richly annotated collection of 1,513 indoor scenes featuring 3D point clouds, RGB images, and camera poses. This shared backbone ensures consistent scene semantics across benchmarks, facilitating multi-task learning. Each benchmark has been aligned to a unified format suitable for instruction-following scenarios. These datasets collectively evaluate a variety of 3D reasoning abilities, ranging from precise localization to compositional understanding.
Metrics.
We use standard evaluation protocols from prior work [22]. For ScanRefer [4], we report thresholded accuracies Acc@0.25 and Acc@0.5, which assess whether the predicted object bounding box has an IoU with the ground truth exceeding 0.25 or 0.5. For Multi3DRefer [58], which involves grounding multiple targets, we use the F1 score at IoU thresholds of 0.25 and 0.5. For the captioning task Scan2Cap [11], we adopt CIDEr@0.5 and BLEU-4@0.5, integrating captioning quality with spatial alignment via IoU. For visual question answering, ScanQA [2] is evaluated using CIDEr and BLEU-4, while SQA3D [34] is evaluated using Exact Match (EM) and its refined variant EM-R as proposed in LEO [23].
4.2 Implementation Details
We extract 100 object proposals per scene using the Mask3D [39] segmentation model. For 3D geometry, we use Uni3D [61] to extract point embeddings $\mathbf{Z}^{v_{p}}_{i}$. For 2D appearance, we adopt DINOv2 [36] to extract 1024-dimensional embeddings per image. We project each object’s 3D mask onto multi-view images and average the cropped DINOv2 embeddings per view. These view-level embeddings are then aggregated using a size-weighted average to obtain the final object-level embeddings $\mathbf{Z}^{v_{f}}_{i}$. For text, we utilize the LLaVA-v1.5-7B [31] model to generate descriptions for each object. These object descriptions are then encoded using the all-mpnet-base-v2 model, a SentenceTransformer [38] based on MPNet [40] that captures rich contextual semantics for downstream fusion, resulting in the embeddings denoted as $\mathbf{Z}^{v_{t}}_{i}$. Each modality is projected to the token space using a three-layer MLP.
We use Vicuna-7B-v1.5 [12] as the LLM and fine-tune it with LoRA [21] with a rank of 16. We train for 3 epochs using a batch size of 32, a base learning rate of $5\times 10^{-6}$, and a cosine annealing schedule. Training is conducted on two 80G NVIDIA A100 GPUs and completes in about 24 hours. We observe that training for 2 epochs yields improved performance on the ScanQA dataset, and we adopt this setting in relevant evaluations.
Method ScanQA SQA3D ScanRefer
BLEU-1 BLEU-4 METEOR ROUGE CIDEr EM EM-R Acc@0.25 Acc@0.5
Expert Models
ScanRefer [4] - - - - - - - 37.3 24.3
ScanQA [2] 30.2 10.1 13.1 33.3 64.9 - - - -
SQA3D [34] - - - - - 46.6 - - -
3DJCG [3] - - - - - - - 49.6 37.3
3D-VLP [27] 30.5 11.1 13.5 34.5 67.0 - - 51.4 39.5
M3DRef-CLIP [58] - - - - - - - 51.9 44.7
3D-VisTA [62] - 13.1 13.9 35.7 72.9 48.5 - 50.6 45.5
ConcreteNet [41] - - - - - - - 50.6 46.5
PQ3D [63] 43.0 - 17.8 - 87.8 47.1 - - 51.2
LLM-based Models
LAMM [52] - 5.8 - - 42.4 - - - 3.4
Chat-3D [44] 29.1 6.4 11.9 28.5 53.2 - - - -
ZSVG3D [54] - - - - - - - 36.4 32.7
3D-LLM [20] 39.3 12.0 14.5 35.7 69.4 - - 30.3 -
LL3DA [8] - 13.5 15.9 37.3 76.8 - - - -
Grounded 3D-LLM [10] - 13.2 - - 75.9 - - 48.6 44.0
LEO [23] - 11.5 16.2 39.3 80.0 50.0 52.4 - -
Scene-LLM [14] 43.6 12.0 16.6 40.0 80.0 54.2 - - -
3DGraphLLM [56] - 12.1 - - 87.6 53.1 - 57.0 51.3
Chat-Scene [22] 43.2 14.3 18.0 41.6 87.7 54.6 57.5 55.5 50.2
Descrip3D (Ours) 44.0 14.5 18.6 43.1 93.7 55.7 58.4 57.2 51.8
Table 1: Performance on ScanQA [2], SQA3D [34], and ScanRefer [4]. “Expert models” are tailored for specific tasks using task-oriented heads, while “LLM-based models” are designed for general instructions and responses. Descrip3D achieves the highest performance across all benchmarks, outperforming prior LLM-based methods by leveraging relational textual descriptions through a dual-level integration for more precise and context-aware reasoning.
4.3 Performance Comparison
To comprehensively evaluate the effectiveness of our proposed relational text modality, we conduct extensive experiments across four representative 3D vision-and-language tasks: 3D visual grounding, question answering, multi-object 3D visual grounding, and scene captioning. These tasks are benchmarked using the five widely adopted datasets mentioned before.
Scene-level Question Answering.
We evaluate our model on ScanQA [2] and SQA3D [34], both of which require comprehensive scene-level understanding and precise object-grounded reasoning. As shown in Table 1, our method achieves the highest scores across all major metrics compared with previous expert models and LLM-based methods, highlighting the strength of our dual-level integrated relational descriptions in enhancing context understanding. Notably, CIDEr, which emphasizes content relevance, shows a significant gain of 5.9 over PQ3D and 6.0 over Chat-Scene, confirming that our model generates more informative and relevant answers. On the SQA3D dataset, our method attains the highest EM and EM-R, again outperforming existing models. The strong results across both benchmarks demonstrate the effectiveness of enriching 3D object representations with object-level textual descriptions. As a general-purpose model, our model consistently outperforms task-specific systems, demonstrating strong versatility without the need for task-dependent architectures. Compared to knowledge-graph-based approaches such as 3DGraphLLM, our model achieves better alignment with language and context, suggesting that lightweight textual descriptions offer a more direct and interpretable semantic grounding. Overall, the results confirm that injecting fine-grained natural language descriptions of objects into the scene representation, through dual-level integration, significantly enhances the LLM’s ability to handle 3D visual question answering tasks.
3D Visual Grounding.
We evaluate 3D visual grounding on the ScanRefer [4] dataset. As shown in Table 1, our method achieves the best performance among all methods. Compared to strong expert models like ConcreteNet and 3D-VisTA, our model surpasses them by 5.3 and 6.3 at the stricter 0.5 IoU threshold, respectively. These expert models are trained with tailored 3D architectures and task-specific objectives, highlighting the strength of our unified and generalizable approach. Among LLM-based methods, our model also clearly outperforms Chat-Scene and 3DGraphLLM. The improvement over Chat-Scene, which uses only 2D and 3D embeddings, validates the great benefit of incorporating relational object-level textual descriptions through a dual-level integration approach.
Multi-Object 3D Visual Grounding.
We evaluate our method on Multi3DRefer [58], a challenging benchmark requiring models to resolve complex multi-object visual grounding within 3D scenes. Unlike traditional single-object grounding, this task demands fine-grained relational reasoning among multiple entities. As shown in Table 2, our method achieves the best performance, outperforming both expert models and LLM-based baselines such as Chat-Scene and Grounded 3D-LLM.
Method Multi3DRefer
F1@0.25 F1@0.5
Expert Models
3DVG-Transformer [59] - 25.5
3DJCG [3] - 26.6
D3Net [5] - 32.2
M3DRef-CLIP [58] 42.8 38.4
PQ3D [63] - 50.1
LLM-based Models
Grounded 3D-LLM [10] 44.7 40.8
Chat-Scene [22] 57.1 52.4
Descrip3D (Ours) 59.4 55.1
Table 2: Performance on Multi3DRefer [58]. Descrip3D achieves the best performance, with notable gains in multi-object reasoning due to its dual-level relational text modeling.
3D Captioning.
We evaluate our method on the Scan2Cap [11] benchmark, which focuses on generating natural language descriptions for 3D objects in complex indoor scenes. As shown in Table 3, our method achieves the highest CIDEr score, indicating that the generated captions are more informative and semantically aligned with the ground truth. While our BLEU-4 score is lower than some other models, this metric emphasizes exact n-gram overlap and can penalize variation in phrasing, particularly in models like ours that incorporate diverse relational language. Importantly, Descrip3D is trained as a unified model across multiple tasks without task-specific tuning, unlike expert models tailored for captioning. Despite this, it still delivers a strong overall performance, suggesting that our relational text descriptions effectively enhance semantic grounding and generalization across diverse scene-language tasks.
Method Scan2Cap
C@0.5 B-4@0.5
Expert Models
Scan2Cap [11] 35.2 22.4
3DJCG [3] 49.5 31.0
3D-VLP [27] 54.9 32.3
3D-VisTA [62] 66.9 34.0
Vote2Cap-DETR++ [9] 67.6 37.1
LLM-based Models
LL3DA [8] 65.2 36.8
LEO [23] 68.4 36.9
Grounded 3D-LLM [10] 70.2 35.0
Chat-Scene [22] 77.1 36.3
Descrip3D (Ours) 77.2 34.5
Table 3: Performance on Scan2Cap [11]. Descrip3D achieves the best C@0.5 performance by leveraging richer object-level relational context.
Qualitative results.
Figure 3 presents qualitative comparisons between our method and Chat-Scene on two representative tasks: 3D question answering (Figure 3(a)) and 3D visual grounding (Figure 3(b)). For question answering, our model demonstrates a stronger ability to understand relational context and resolve spatial references. In comparison, Chat-Scene demonstrably struggles more with object relationships. For visual grounding, our model more accurately identifies target objects based on dual-level integrated complex natural language descriptions that involve appearance and relative position. In contrast, Chat-Scene frequently fails to disambiguate between visually or semantically similar candidates. These results highlight the benefit of injecting detailed object-level descriptions into the language model through a dual-level approach, enabling more precise and interpretable multi-modal reasoning.
(a) Qualitative Comparison of Question Answering.
(b) Qualitative Comparison of Grounding.
Figure 3: Qualitative comparison of 3D scene understanding tasks. Descrip3D outperforms Chat-Scene, especially in cases involving complex spatial grounding or multi-object reasoning, due to its use of a dual-level integrated relational textual descriptions that enhance contextual understanding.
In summary, our method achieves strong and consistent results across the five benchmarks, demonstrating its effectiveness in different tasks.
4.4 Ablation Studies
To evaluate the impact of our proposed object-level text descriptions and the dual-level integration strategies, we conduct a series of ablation experiments across key components of our model.
Effect of Object Reference Style in Prompt-Level Description.
We examine how the choice of object reference expression in prompt-level text injection affects performance. As shown in Table 4, using raw object names can lead to suboptimal performance due to ambiguity when multiple instances of the same category are present. Appending object IDs to names may confuse the model further, failing to improve performance and potentially introducing redundancy. In contrast, using object IDs alone yields the best results across all benchmarks. This strategy eliminates ambiguity and ensures precise alignment between the referenced objects in the query and their corresponding descriptions.
Reference Style ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
Object Name Only 51.6 54.5 75.6 91.4 55.1
Object Name + ID 51.1 53.9 74.1 91.7 54.2
Object ID Only (Ours) 51.8 55.1 77.2 93.7 55.7
Table 4: Ablation study on object reference style in prompt-level text injection. Using object IDs alone yields the best performance across all benchmarks, as it avoids ambiguity and improves alignment between queries and object descriptions.
Effect of Object Descriptions and Fusion Location.
We study the impact of incorporating object-level textual descriptions at different points in the model. As shown in Table 5, injecting textual descriptions solely as embeddings improves grounding and question answering by enriching object semantics. Injecting them into the prompt also boosts performance on tasks like ScanQA and Scan2Cap. Combining both yields the best results on most datasets, highlighting their complementarity. However, on Multi3DRefer, this strategy slightly underperforms, possibly due to redundant signals in complex scenes. Overall, the dual-level integration enhances both spatial grounding and semantic reasoning.
Embedding Prompt ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
50.2 52.4 77.1 87.7 54.6
51.5 55.4 74.6 89.3 54.5
51.7 55.5 75.3 92.3 53.6
51.8 55.1 77.2 93.7 55.7
Table 5: Ablation study on text description and injection strategy. Combining embedding-level and prompt-level injection consistently leads to the best overall performance, demonstrating the complementary benefits of dual-level relational text integration.
5 Conclusion and Future Work
We propose Descrip3D, a simple yet powerful framework for 3D scene understanding that explicitly models inter-object relationships using object-level textual descriptions. By integrating these relational cues through a dual-level strategy, embedding fusion and prompt injection, our method enables more effective multi-object reasoning.
Experiments on five benchmarks show that Descrip3D consistently outperforms expert and LLM-based models, with ablations confirming the importance of dual-level integration. Our results highlight the strength of language as a medium for structured scene representation. As future work, we plan to explore end-to-end training that jointly learns object description generation and extend our approach to dynamic or outdoor environments.
Reference
  • [1] Panos Achlioptas, Ahmed Abdelreheem, Fei Xia, Mohamed Elhoseiny, and Leonidas Guibas. Referit3d: Neural listeners for fine-grained 3d object identification in real-world scenes. In European conference on computer vision, pages 422–440. Springer, 2020.
  • [2] Daichi Azuma, Taiki Miyanishi, Shuhei Kurita, and Motoaki Kawanabe. Scanqa: 3d question answering for spatial scene understanding. In proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19129–19139, 2022.
  • [3] Daigang Cai, Lichen Zhao, Jing Zhang, Lu Sheng, and Dong Xu. 3djcg: A unified framework for joint dense captioning and visual grounding on 3d point clouds. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 16464–16473, 2022.
  • [4] Dave Zhenyu Chen, Angel X Chang, and Matthias Nießner. Scanrefer: 3d object localization in rgb-d scans using natural language. In European conference on computer vision, pages 202–221. Springer, 2020.
  • [5] Dave Zhenyu Chen, Qirui Wu, Matthias Nießner, and Angel X. Chang. D3net: A speaker-listener architecture for semi-supervised dense captioning and visual grounding in rgb-d scans, 2021.
  • [6] Jiaming Chen, Weixin Luo, Xiaolin Wei, Lin Ma, and Wei Zhang. Ham: Hierarchical attention model with high performance for 3d visual grounding. arXiv preprint arXiv:2210.12513, 2(3), 2022.
  • [7] Shizhe Chen, Pierre-Louis Guhur, Makarand Tapaswi, Cordelia Schmid, and Ivan Laptev. Language conditioned spatial relation reasoning for 3d object grounding. Advances in neural information processing systems, 35:20522–20535, 2022.
  • [8] Sijin Chen, Xin Chen, Chi Zhang, Mingsheng Li, Gang Yu, Hao Fei, Hongyuan Zhu, Jiayuan Fan, and Tao Chen. Ll3da: Visual interactive instruction tuning for omni-3d understanding reasoning and planning. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 26428–26438, 2024.
  • [9] Sijin Chen, Hongyuan Zhu, Mingsheng Li, Xin Chen, Peng Guo, Yinjie Lei, Gang Yu, Taihao Li, and Tao Chen. Vote2cap-detr++: Decoupling localization and describing for end-to-end 3d dense captioning. IEEE Transactions on Pattern Analysis and Machine Intelligence, 46(11):7331–7347, 2024.
  • [10] Yilun Chen, Shuai Yang, Haifeng Huang, Tai Wang, Runsen Xu, Ruiyuan Lyu, Dahua Lin, and Jiangmiao Pang. Grounded 3d-llm with referent tokens. arXiv preprint arXiv:2405.10370, 2024.
  • [11] Zhenyu Chen, Ali Gholami, Matthias Nießner, and Angel X Chang. Scan2cap: Context-aware dense captioning in rgb-d scans. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 3193–3203, 2021.
  • [12] Wei-Lin Chiang, Zhuohan Li, Ziqing Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E Gonzalez, et al. Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality. See https://vicuna. lmsys. org (accessed 14 April 2023), 2(3):6, 2023.
  • [13] Angela Dai, Angel X Chang, Manolis Savva, Maciej Halber, Thomas Funkhouser, and Matthias Nießner. Scannet: Richly-annotated 3d reconstructions of indoor scenes. In Proceedings of the IEEE conference on computer vision and pattern recognition, pages 5828–5839, 2017.
  • [14] Rao Fu, Jingyu Liu, Xilun Chen, Yixin Nie, and Wenhan Xiong. Scene-llm: Extending language model for 3d visual understanding and reasoning. arXiv preprint arXiv:2403.11401, 2024.
  • [15] Qiao Gu, Ali Kuwajerwala, Sacha Morin, Krishna Murthy Jatavallabhula, Bipasha Sen, Aditya Agarwal, Corban Rivera, William Paul, Kirsty Ellis, Rama Chellappa, et al. Conceptgraphs: Open-vocabulary 3d scene graphs for perception and planning. In 2024 IEEE International Conference on Robotics and Automation (ICRA), pages 5021–5028. IEEE, 2024.
  • [16] Ziyu Guo, Renrui Zhang, Xiangyang Zhu, Yiwen Tang, Xianzheng Ma, Jiaming Han, Kexin Chen, Peng Gao, Xianzhi Li, Hongsheng Li, et al. Point-bind & point-llm: Aligning point cloud with multi-modality for 3d understanding, generation, and instruction following. arXiv preprint arXiv:2309.00615, 2023.
  • [17] Muhammad Usman Hadi, Rizwan Qureshi, Abbas Shah, Muhammad Irfan, Anas Zafar, Muhammad Bilal Shaikh, Naveed Akhtar, Jia Wu, Seyedali Mirjalili, et al. A survey on large language models: Applications, challenges, limitations, and practical usage. Authorea Preprints, 2023.
  • [18] Jiaming Han, Renrui Zhang, Wenqi Shao, Peng Gao, Peng Xu, Han Xiao, Kaipeng Zhang, Chris Liu, Song Wen, Ziyu Guo, et al. Imagebind-llm: Multi-modality instruction tuning. arXiv preprint arXiv:2309.03905, 2023.
  • [19] Deepti Hegde, Jeya Maria Jose Valanarasu, and Vishal Patel. Clip goes 3d: Leveraging prompt tuning for language grounded 3d recognition. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2028–2038, 2023.
  • [20] Yining Hong, Haoyu Zhen, Peihao Chen, Shuhong Zheng, Yilun Du, Zhenfang Chen, and Chuang Gan. 3d-llm: Injecting the 3d world into large language models. Advances in Neural Information Processing Systems, 36:20482–20494, 2023.
  • [21] Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, Weizhu Chen, et al. Lora: Low-rank adaptation of large language models. ICLR, 1(2):3, 2022.
  • [22] Haifeng Huang, Yilun Chen, Zehan Wang, Rongjie Huang, Runsen Xu, Tai Wang, Luping Liu, Xize Cheng, Yang Zhao, Jiangmiao Pang, et al. Chat-scene: Bridging 3d scene and large language models with object identifiers. Advances in Neural Information Processing Systems, 37:113991–114017, 2024.
  • [23] Jiangyong Huang, Silong Yong, Xiaojian Ma, Xiongkun Linghu, Puhao Li, Yan Wang, Qing Li, Song-Chun Zhu, Baoxiong Jia, and Siyuan Huang. An embodied generalist agent in 3d world. arXiv preprint arXiv:2311.12871, 2023.
  • [24] Pin-Hao Huang, Han-Hung Lee, Hwann-Tzong Chen, and Tyng-Luh Liu. Text-guided graph neural networks for referring 3d instance segmentation. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 35, pages 1610–1618, 2021.
  • [25] Shijia Huang, Yilun Chen, Jiaya Jia, and Liwei Wang. Multi-view transformer for 3d visual grounding. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 15524–15533, 2022.
  • [26] Ayush Jain, Nikolaos Gkanatsios, Ishita Mediratta, and Katerina Fragkiadaki. Bottom up top down detection transformers for language grounding in images and point clouds. In European Conference on Computer Vision, pages 417–433. Springer, 2022.
  • [27] Zhao Jin, Munawar Hayat, Yuwei Yang, Yulan Guo, and Yinjie Lei. Context-aware alignment and mutual masking for 3d-language pre-training. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 10984–10994, 2023.
  • [28] Xinyi Li, Sai Wang, Siqi Zeng, Yu Wu, and Yi Yang. A survey on llm-based multi-agent systems: workflow, infrastructure, and challenges. Vicinagearth, 1(1):9, 2024.
  • [29] Zechuan Li, Hongshan Yu, Yihao Ding, Yan Li, Yong He, and Naveed Akhtar. Embodied intelligence for 3d understanding: A survey on 3d scene question answering. arXiv preprint arXiv:2502.00342, 2025.
  • [30] Daizong Liu, Yang Liu, Wencan Huang, and Wei Hu. A survey on text-guided 3d visual grounding: Elements, recent advances, and future directions. arXiv preprint arXiv:2406.05785, 2024.
  • [31] Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. Visual instruction tuning. In NeurIPS, 2023.
  • [32] Junyu Luo, Jiahui Fu, Xianghao Kong, Chen Gao, Haibing Ren, Hao Shen, Huaxia Xia, and Si Liu. 3d-sps: Single-stage 3d visual grounding via referred point progressive selection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 16454–16463, 2022.
  • [33] Xianzheng Ma, Yash Bhalgat, Brandon Smart, Shuai Chen, Xinghui Li, Jian Ding, Jindong Gu, Dave Zhenyu Chen, Songyou Peng, Jia-Wang Bian, et al. When llms step into the 3d world: A survey and meta-analysis of 3d tasks via multi-modal large language models. arXiv preprint arXiv:2405.10255, 2024.
  • [34] Xiaojian Ma, Silong Yong, Zilong Zheng, Qing Li, Yitao Liang, Song-Chun Zhu, and Siyuan Huang. Sqa3d: Situated question answering in 3d scenes. arXiv preprint arXiv:2210.07474, 2022.
  • [35] Shervin Minaee, Tomas Mikolov, Narjes Nikzad, Meysam Chenaghlu, Richard Socher, Xavier Amatriain, and Jianfeng Gao. Large language models: A survey. arXiv preprint arXiv:2402.06196, 2024.
  • [36] Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al. Dinov2: Learning robust visual features without supervision. arXiv preprint arXiv:2304.07193, 2023.
  • [37] Zhangyang Qi, Zhixiong Zhang, Ye Fang, Jiaqi Wang, and Hengshuang Zhao. Gpt4scene: Understand 3d scenes from videos with vision-language models. arXiv preprint arXiv:2501.01428, 2025.
  • [38] Nils Reimers and Iryna Gurevych. Sentence-bert: Sentence embeddings using siamese bert-networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, 11 2019.
  • [39] Jonas Schult, Francis Engelmann, Alexander Hermans, Or Litany, Siyu Tang, and Bastian Leibe. Mask3d: Mask transformer for 3d semantic instance segmentation. arXiv preprint arXiv:2210.03105, 2022.
  • [40] Kaitao Song, Xu Tan, Tao Qin, Jianfeng Lu, and Tie-Yan Liu. Mpnet: Masked and permuted pre-training for language understanding. Advances in neural information processing systems, 33:16857–16867, 2020.
  • [41] Ozan Unal, Christos Sakaridis, Suman Saha, Fisher Yu, and Luc Van Gool. Three ways to improve verbo-visual fusion for dense 3d visual grounding. arXiv preprint arXiv:2309.04561, 2:15, 2023.
  • [42] Yun-Cheng Wang, Jintang Xue, Chengwei Wei, and C-C Jay Kuo. An overview on generative ai at scale with edge–cloud computing. IEEE Open Journal of the Communications Society, 4:2952–2971, 2023.
  • [43] Zehan Wang, Haifeng Huang, Yang Zhao, Linjun Li, Xize Cheng, Yichen Zhu, Aoxiong Yin, and Zhou Zhao. 3drp-net: 3d relative position-aware network for 3d visual grounding. arXiv preprint arXiv:2307.13363, 2023.
  • [44] Zehan Wang, Haifeng Huang, Yang Zhao, Ziang Zhang, and Zhou Zhao. Chat-3d: Data-efficiently tuning large language model for universal dialogue of 3d scenes. arXiv preprint arXiv:2308.08769, 2023.
  • [45] Abdelrhman Werby, Chenguang Huang, Martin Büchner, Abhinav Valada, and Wolfram Burgard. Hierarchical open-vocabulary 3d scene graphs for language-grounded robot navigation. In First Workshop on Vision-Language Models for Navigation and Manipulation at ICRA 2024, 2024.
  • [46] Tung-Yu Wu, Sheng-Yu Huang, and Yu-Chiang Frank Wang. Dora: 3d visual grounding with order-aware referring. CoRR, 2024.
  • [47] Yanmin Wu, Xinhua Cheng, Renrui Zhang, Zesen Cheng, and Jian Zhang. Eda: Explicit text-decoupling and dense alignment for 3d visual grounding. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19231–19242, 2023.
  • [48] Runsen Xu, Xiaolong Wang, Tai Wang, Yilun Chen, Jiangmiao Pang, and Dahua Lin. Pointllm: Empowering large language models to understand point clouds. In European Conference on Computer Vision, pages 131–147. Springer, 2024.
  • [49] Jintang Xue, Yun-Cheng Wang, Chengwei Wei, Xiaofeng Liu, Jonghye Woo, C-C Jay Kuo, et al. Bias and fairness in chatbots: An overview. APSIPA Transactions on Signal and Information Processing, 13(2), 2024.
  • [50] Le Xue, Mingfei Gao, Chen Xing, Roberto Martín-Martín, Jiajun Wu, Caiming Xiong, Ran Xu, Juan Carlos Niebles, and Silvio Savarese. Ulip: Learning a unified representation of language, images, and point clouds for 3d understanding. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 1179–1189, 2023.
  • [51] Zhengyuan Yang, Songyang Zhang, Liwei Wang, and Jiebo Luo. Sat: 2d semantics assisted training for 3d visual grounding. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1856–1866, 2021.
  • [52] Zhenfei Yin, Jiong Wang, Jianjian Cao, Zhelun Shi, Dingning Liu, Mukai Li, Xiaoshui Huang, Zhiyong Wang, Lu Sheng, Lei Bai, et al. Lamm: Language-assisted multi-modal instruction-tuning dataset, framework, and benchmark. Advances in Neural Information Processing Systems, 36:26650–26685, 2023.
  • [53] Ting Yu, Xiaojun Lin, Shuhui Wang, Weiguo Sheng, Qingming Huang, and Jun Yu. A comprehensive survey of 3d dense captioning: Localizing and describing objects in 3d scenes. IEEE Transactions on Circuits and Systems for Video Technology, 34(3):1322–1338, 2023.
  • [54] Zhihao Yuan, Jinke Ren, Chun-Mei Feng, Hengshuang Zhao, Shuguang Cui, and Zhen Li. Visual programming for zero-shot open-vocabulary 3d visual grounding. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 20623–20633, 2024.
  • [55] Zhihao Yuan, Xu Yan, Yinghong Liao, Ruimao Zhang, Sheng Wang, Zhen Li, and Shuguang Cui. Instancerefer: Cooperative holistic understanding for visual grounding on point clouds through instance multi-level contextual referring. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1791–1800, 2021.
  • [56] Tatiana Zemskova and Dmitry Yudin. 3dgraphllm: Combining semantic graphs and large language models for 3d scene understanding. arXiv preprint arXiv:2412.18450, 2024.
  • [57] Jirong Zha, Yuxuan Fan, Xiao Yang, Chen Gao, and Xinlei Chen. How to enable llm with 3d capacity? a survey of spatial reasoning in llm. arXiv preprint arXiv:2504.05786, 2025.
  • [58] Yiming Zhang, ZeMing Gong, and Angel X Chang. Multi3drefer: Grounding text description to multiple 3d objects. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 15225–15236, 2023.
  • [59] Lichen Zhao, Daigang Cai, Lu Sheng, and Dong Xu. 3dvg-transformer: Relation modeling for visual grounding on point clouds. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2928–2937, 2021.
  • [60] Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. A survey of large language models. arXiv preprint arXiv:2303.18223, 1(2), 2023.
  • [61] Junsheng Zhou, Jinsheng Wang, Baorui Ma, Yu-Shen Liu, Tiejun Huang, and Xinlong Wang. Uni3d: Exploring unified 3d representation at scale. arXiv preprint arXiv:2310.06773, 2023.
  • [62] Ziyu Zhu, Xiaojian Ma, Yixin Chen, Zhidong Deng, Siyuan Huang, and Qing Li. 3d-vista: Pre-trained transformer for 3d vision and text alignment. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2911–2921, 2023.
  • [63] Ziyu Zhu, Zhuofan Zhang, Xiaojian Ma, Xuesong Niu, Yixin Chen, Baoxiong Jia, Zhidong Deng, Siyuan Huang, and Qing Li. Unifying 3d vision-language understanding via promptable queries. In European Conference on Computer Vision, pages 188–206. Springer, 2024.
Appendix A Examples of Generated Descriptions
Figure 4: Qualitative Examples of Object-level Relational Descriptions Generated from Multi-view Images. The upper left part displays the image with object names, and the lower left part shows the bounding boxes of the objects.
We present qualitative examples of our object-level textual descriptions in Figure 4. Starting from detected object proposals and their corresponding multi-view images, we overlay the object names at the center of the projection areas in the image, as illustrated in the upper left part of the example. We then generate relational descriptions using a vision-language model. Key objects, typically those centrally positioned in the scene, are selected as query anchors. For each key object, we prompt the model to describe its spatial relationships with all other detected objects, resulting in detailed, contextually grounded descriptions. The prompt used is: “Describe clearly and briefly the relationships between the <Key Object>in the scene and nearby objects (<Other Object 1>, <Other Object 2>, …, <Other Object n>). Do not describe objects you cannot see.” For example, the objects in the image are a desk, two curtains, a window, a cabinet, and a table. There are two curtains, but only the one on the right is considered a key object because the other is positioned at the edge of the image. The chosen curtain is described as covering the window and situated near the table, the cabinet, and the desk. These relational descriptions offer interpretable summaries of local neighborhoods and equip downstream models with structured scene understanding for improved reasoning.
Appendix B Ablation Study on Object Labels in Description Generation
To examine the impact of explicitly overlaying object category names during relational description generation, we conduct an ablation study comparing two variants: one where multi-view images include projected object labels (ours), and one without. As shown in Table 6, incorporating object labels consistently improves performance across all five benchmarks. The improvement is particularly notable in Scan2Cap and SQA3D, where more precise object references in the descriptions likely benefit caption generation and question answering. These results confirm that providing explicit category labels helps the vision-language model better ground each object and generate more informative relational descriptions.
Multi-view Image Input ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
Without Object Labels 51.5 54.8 75.6 93.5 54.6
With Object Labels (Ours) 51.8 55.1 77.2 93.7 55.7
Table 6: Ablation study on the effect of overlaying object category labels in multi-view images during relational description generation. Adding object labels leads to consistent performance improvements across all benchmarks, demonstrating their importance in guiding the vision-language model toward accurate grounding.
Appendix C Additional Quantitative Results
We evaluate our method using the standard metrics established in the original papers for each 3D scene-language dataset. To thoroughly assess the effectiveness of our approach, we perform extensive comparisons against a diverse set of baselines across multiple benchmarks. To complement the main results, we report additional evaluation metrics on the same datasets (ScanRefer, Multi3DRefer, and ScanQA) used in the main paper. The results, summarized in Table 8 (ScanRefer), Table 9 (Multi3DRefer), and Table 10 (ScanQA), show our method consistently outperforms prior approaches across grounding and question answering tasks. On ScanRefer, Descrip3D achieves the highest overall accuracy. On Multi3DRefer, it leads in almost all grounding settings, with the best overall F1 scores. On ScanQA, it outperforms baselines in nearly all language metrics, including ROUGE-L, METEOR, and CIDEr. These results confirm the effectiveness of incorporating object-level textual descriptions through dual-level integration for 3D vision-language tasks.
Appendix D Prompt Template
[…]
User: What is the ID of the object that matches the description "this is a long table. it is surrounded by chairs"?
Assistant: <OBJ023>
Table 7: Prompt template used during training and evaluation.
Method Venue Unique Multiple Overall
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [4] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
TGNN [24] AAAI21 68.61 56.80 29.84 23.18 37.37 29.70
SAT [51] ICCV21 73.21 50.83 37.64 25.16 44.54 30.14
InstanceRefer [55] ICCV21 75.72 64.66 29.41 22.99 38.40 31.08
3DVG-Transformer [59] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
MVT [25] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
3D-SPS [32] CVPR22 84.12 66.72 40.32 29.82 48.82 36.98
ViL3DRel [7] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
3DJCG [3] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
D3Net [5] ECCV22 72.04 30.05 37.87
BUTD-DETR [26] ECCV22 84.2 66.3 46.6 35.1 52.2 39.8
HAM [6] ArXiv22 79.24 67.86 41.46 34.03 48.79 40.60
3DRP-Net [43] EMNLP23 83.13 67.74 42.14 31.95 50.10 38.90
3D-VLP [27] CVPR23 84.23 64.61 43.51 33.41 51.41 39.46
EDA [47] CVPR23 85.76 68.57 49.13 37.64 54.59 42.26
M3DRef-CLIP [58] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [62] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
ConcreteNet [41] ECCV24 86.40 82.05 42.41 38.39 50.61 46.53
DORa [46] ArXiv24 52.80 44.80
Chat-Scene [22] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
Descrip3D (Ours) 90.79 83.23 49.62 44.72 57.24 51.84
Table 8: Performance comparison on the validation set of ScanRefer [4].
Method Venue ZT w/o D ZT w/ D ST w/o D ST w/ D MT ALL
F1 F1 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5
3DVG-Trans+ [59] ICCV21 87.1 45.8 16.7 26.5 25.5
D3Net (Grounding) [5] ECCV22 81.6 32.5 23.3 35.0 32.2
3DJCG (Grounding) [3] CVPR22 94.1 66.9 16.7 26.2 26.6
M3DRef-CLIP [58] ICCV23 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
Chat-Scene [22] NeurIPS24 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
Descrip3D (Ours) 92.0 70.4 83.1 75.9 51.4 47.4 49.2 45.2 59.4 55.1
Table 9: Performance comparison on the validation set of Multi3DRefer [58].
Method Venue EM@1 B-1 B-2 B-3 B-4 ROUGE-L METEOR CIDEr
ScanQA [2] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
3D-VLP [27] CVPR22 21.65 30.53 21.33 16.67 11.15 34.51 13.53 66.97
3D-LLM [20] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
LL3DA [8] CVPR24 13.53 37.31 15.88 76.79
LEO [23] ICML24 11.5 39.3 16.2 80.0
Scene-LLM [14] WACV25 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
Chat-Scene [22] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
Descrip3D (Ours) 22.67 44.36 30.51 22.08 15.70 43.01 19.06 93.71
Table 10: Performance comparison on the validation set of ScanQA [2].

### v2

@@
 When evaluated on five benchmark datasets, including ScanRefer, Multi3DRefer, ScanQA, SQA3D, and Scan2Cap, Descrip3D consistently outperforms strong baseline models, demonstrating the effectiveness of language-guided relational representation for understanding complex indoor scenes.
+Our code and data are publicly available at https://github.com/jintangxue/Descrip3D.
 1 Introduction
+Figure 1: An example of injecting object-level text descriptions during conversation. Providing these descriptions significantly improves the model’s accuracy and reasoning performance.
 Recent advances in large language models (LLMs) have significantly transformed human-computer interaction by equipping machines with unprecedented capabilities in language understanding, reasoning, and open-ended dialogue.
-These models have demonstrated remarkable success in a wide range of domains, including information retrieval, code generation, and multi-modal learning [17, 35, 60, 28, 49, 42].
-Building on this progress, researchers have begun to extend LLMs to 3D scene understanding [33, 57, 37], aiming to empower models with the ability to interpret and reason about complex visual and spatial contexts within real-world environments.
-In particular, indoor scenes present unique challenges for 3D scene-language modeling, including dense layouts, ambiguous viewpoints, and complex spatial relationships.
-Earlier methods [4, 11] were typically designed for specific tasks such as grounding or captioning.
-In contrast, recent efforts [20, 22] explore 3D multimodal large language models (MLLMs) that integrate language with point clouds and multiview images to jointly support a wide range of tasks, such as 3D visual grounding, dense captioning, and question answering, using a single trained model.
-Models like Chat-Scene [22] adopt object-centric pipelines by associating 2D/3D embeddings with object identifiers and feeding them into the language model.
-However, these approaches still show limited performance on tasks requiring spatial or semantic reasoning across multiple objects, primarily because they lack explicit modeling of inter-object relationships and underutilize the language understanding capabilities of LLMs.
-Although recent graph-based methods [56] attempt to address this, they introduce additional training complexity and still fall short in performance.
-Figure 1: An example of injecting object-level text descriptions during conversation. Providing these descriptions significantly improves the model’s accuracy and reasoning performance.
+LLMs have achieved strong results in domains such as retrieval, code, and multi-modal learning [hadi2023survey, minaee2024large, zhao2023survey, li2024survey, xue2024bias, wang2023overview].
+Building on this progress, researchers have begun to extend LLMs to 3D scene understanding [ma2024llms, zha2025enable, qi2025gpt4scene], aiming to empower models with the ability to interpret and reason about complex visual and spatial contexts within real-world environments.
+In particular, indoor scenes are dense and ambiguous.
+While early models [chen2020scanrefer, chen2021scan2cap] targeted single tasks and recent LLMs [hong20233d, huang2024chat] integrate 2D/3D features, none explicitly model inter-object relations.
 As illustrated in Figure 1, we address this limitation by introducing a novel text-based relational modality into the 3D scene language pipeline.
@@
 This dual-level integration, at both the representation and the language interface levels, allows the model to leverage the prior knowledge of the LLM about the spatial and functional relations learned from the language.
-Crucially, this approach does not introduce architectural changes, does not introduce task-specific heads, and does not require additional human annotation.
-It is modular, generalizable, and easy to integrate into existing object-centric pipelines.
+Our approach requires no architectural changes, task-specific heads, or extra annotations, making it modular and easy to integrate.
 Empirically, our method yields consistent gains across five standard 3D scene language benchmarks.
-ScanRefer [4], Multi3DRefer [58], Scan2Cap [11], ScanQA [2], and SQA3D [34].
+ScanRefer [chen2020scanrefer], Multi3DRefer [zhang2023multi3drefer], Scan2Cap [chen2021scan2cap], ScanQA [azuma2022scanqa], and SQA3D [ma2022sqa3d].
 Improvements are especially pronounced in tasks involving complex grounding or multi-object reasoning, highlighting the benefit of explicit text-based relational cues in enhancing 3D scene understanding.
+Here, multi-object reasoning does not refer only to Multi3DRefer’s multi-target grounding protocol.
+We use the term more generally to mean queries whose solution depends on object–object relations (e.g., left of, behind, next to, supporting, inside), regardless of whether the benchmark requires multiple boxes.
 Our contributions are as follows.
@@
 2 Related Work
-2.1 3D Vision-Language Understanding
-3D vision-language understanding combines 3D spatial perception with natural language to interpret and interact with complex environments.
-Earlier work focused on single-task models for visual grounding [30], dense captioning [53], and question answering [29].
-For example, models such as ScanRefer [4], ReferIt3D [1], and Multi3DRefer [58] focus on grounding by locating target objects based on descriptions of natural language.
-In terms of dense captioning, models such as the 3DVG-Transformer [59] and Vote2Cap-DETR++ [9] work to both localize and describe objects within a 3D environment.
-Furthermore, question-answer models like ScanQA [2] and SQA3D [34] extend the task space by incorporating a broader understanding of the scene to answer general or context-dependent questions.
-Recent research is increasingly focusing on multi-task frameworks that unify these objectives to enhance representation sharing and robust performance.
-For example, methods such as 3DJCG [3] and D3Net [5] unify 3D visual grounding and dense captioning through shared encoders and task-specific heads, leveraging complementary supervision across tasks to improve performance.
-Likewise, 3D-VisTA [62] and 3D-VLP [27] employ large-scale pre-training to learn generalizable 3D vision-language representations, enhancing performance across diverse downstream tasks.
-However, their reliance on task-specific heads and visual embeddings treats language as auxiliary, lacking explicit modeling of semantic object relationships.
-2.2 Multi-modal Large Language Models
-LLMs have demonstrated impressive capabilities in reasoning, generalization, and multi-turn dialogue, sparking interest in extending their input space to include 3D data [33].
-In the 3D vision language domain, models such as PointLLM [48], ImageBind-LLM [18], Point-Bind [16], Ulip [50], and CG3D [19] extend LLMs to handle 3D data by aligning the embeddings of the point cloud with the language embedding space.
-These methods enable object-level understanding through joint representations of geometric and linguistic information, facilitating tasks such as 3D object captioning and question answering.
-However, they primarily operate on isolated object input and lack mechanisms for encoding explicit relationships between objects, which are essential for capturing scene-level semantics.
-Multimodal large language models (MLLMs) have recently been extended to 3D scene understanding to support open-ended tasks such as question answering.
-Approaches such as 3D-LLM [20] map the embeddings of point clouds into the language space using position embeddings, allowing basic multimodal reasoning but lacking fine-grained localization and interactive grounding.
-Chat-3D [44] enhances the 3D dialogue by using object-centric prompts and a three-stage training scheme to align the characteristics and relations of objects with language models.
-However, its architecture restricts the model’s attention to individually selected objects, limiting its ability to reason in broader contexts of the scene.
-Chat-Scene [22] further advanced this line by incorporating object identifiers and aligning them with 2D and 3D visual embeddings, enabling strong performance on scene-level tasks such as question answering and visual grounding.
-Yet, they still lack explicit mechanisms for modeling inter-object relationships, limiting relational and spatial reasoning.
+2.1 LLMs for 3D Scene Understanding
+Research in 3D vision-language has progressed from task-specific solutions to unified and LLM-driven frameworks.
+Earlier work focused on single-task models for visual grounding [liu2024survey], dense captioning [yu2023comprehensive], and question answering [li2025embodied].
+For example, ScanRefer [chen2020scanrefer], ReferIt3D [achlioptas2020referit3d], and Multi3DRefer [zhang2023multi3drefer] tackled visual grounding by linking natural language to 3D objects.
+For dense captioning, 3DVG-Transformer [zhao20213dvg] and Vote2Cap-DETR++ [chen2024vote2cap] localized and described objects, while ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d] broadened the scope to open-ended scene-level QA.
+Building on these, multi-task frameworks such as 3DJCG [cai20223djcg] and D3Net [chen2022d3net] unified grounding and captioning with shared encoders, and large-scale pretraining approaches like 3D-VisTA [zhu20233d] and 3D-VLP [jin2023context] established transferable 3D vision-language representations.
+Yet most treat language as auxiliary, limiting explicit relational modeling.
+The emergence of large language models (LLMs) has introduced stronger reasoning and conversational capabilities, motivating their integration with 3D data.
+Approaches such as PointLLM [xu2024pointllm], ImageBind-LLM [han2023imagebind], Point-Bind [guo2023point], Ulip [xue2023ulip], and CG3D [hegde2023clip] align point cloud embeddings with text, enabling object-level captioning and QA.
+However, these mainly operate on isolated objects, lacking mechanisms for modeling interactions within complex scenes.
+To overcome this, multimodal LLMs (MLLMs) have been extended to scene-level reasoning.
+3D-LLM [hong20233d] projects 3D embeddings into language space for basic multimodal reasoning, while Chat-3D [wang2023chat] introduces object-centric prompts and staged training to align 3D objects with dialogue.
+More recently, Chat-Scene [huang2024chat] incorporates object identifiers and fuses them with 2D and 3D embeddings, achieving strong results in QA and grounding benchmarks.
+Yet, even these advances fall short of explicitly encoding inter-object relationships, leaving relational and spatial reasoning an open challenge for LLM-based 3D scene understanding.
 2.2 Explicit Object Relationship Representation
 Although prior models lack explicit mechanisms for encoding inter-object relationships, recent efforts have begun to address this limitation through graph-based or language-based modeling.
-ConceptGraph [15] constructs scene-level concept graphs by aligning 3D object embeddings with textual descriptions using vision-language models, enabling semantic relationship modeling and open-vocabulary generalization.
-HOV-SG [45] advances this direction by generating holistic object-scene graphs from 3D scans, capturing spatial and semantic relations across the scene.
-However, such graph-based representations often require long token sequences and suffer from high inference costs in large-scale scenes.
-To address this, 3DGraphLLM [56] incorporates knowledge graph embeddings into object representations, encoding pairwise relationships to support relational reasoning.
+ConceptGraph [gu2024conceptgraphs] constructs scene-level concept graphs by aligning 3D object embeddings with textual descriptions using vision-language models, enabling semantic relationship modeling and open-vocabulary generalization.
+HOV-SG [werby2024hierarchical] advances this direction by generating holistic object-scene graphs from 3D scans, capturing spatial and semantic relations across the scene.
+However, graph-based representations require long sequences and incur high inference costs.
+To address this, 3DGraphLLM [zemskova20243dgraphllm] incorporates knowledge graph embeddings into object representations, encoding pairwise relationships to support relational reasoning.
 Yet, this approach introduces significant combinatorial overhead by enumerating object-object triplets and suffers from error propagation in graph construction.
 Notably, the addition of explicit graph structures in 3DGraphLLM leads to a performance drop on downstream question answering tasks, suggesting that current graph-based formulations may hinder rather than help relational reasoning in practice.
-Scene-LLM [14] instead leverages LLMs with scene-level prompts to generate global summaries and captions, but lacks fine-grained object-level relational modeling.
+Scene-LLM [fu2024scene] instead leverages LLMs with scene-level prompts to generate global summaries and captions, but lacks fine-grained object-level relational modeling.
 Thus, explicitly and efficiently representing object relationships in 3D scenes remains an open and challenging problem.
+2.3 3D Dataset Construction via 2D VLMs
+Recent works leverage 2D vision-language models to expand 3D–text resources.
+Cap3D [luo2023scalable] generates captions for millions of Objaverse objects by aggregating multi-view 2D renderings with pretrained captioning and alignment models.
+LiDAR-LLM [yang2025lidar] reformulates outdoor LiDAR understanding as a language modeling task, producing hundreds of thousands of LiDAR–text pairs through a staged pipeline.
+These efforts primarily focus on large-scale data creation to support 3D–language research.
+In contrast, Descrip3D does not rely on new annotations; instead, it injects object-level descriptions directly into the model pipeline to strengthen relational reasoning across existing benchmarks.
 3 Method
@@
 This dual-level integration significantly improves the ability of LLM to reason about object relationships, supporting downstream tasks such as grounding, captioning, and question answering with high accuracy and interpretability.
-An overview of our pipeline is shown in Figure 2.
+An overview of our pipeline is shown in fig.˜2.
 Figure 2: Overall model architecture.
@@
 3.2 Object-Centric Scene Representation
-We decompose each scene into discrete object proposals using Mask3D [39] first, producing a set of segmented object point clouds $\{\mathbf{P}_{1},...,\mathbf{P}_{n}\}$.
+We decompose each scene into discrete object proposals using Mask3D [schult2022mask3d] first, producing a set of segmented object point clouds $\{\mathbf{P}_{1},...,\mathbf{P}_{n}\}$.
 Each object $\mathbf{P}_{i}\in\mathbb{R}^{m_{i}\times 6}$ includes XYZ coordinates and RGB values.
@@
 Each resulting object is treated as a fundamental unit for understanding the scene.
-To support unambiguous reference and interaction, we adopt the design introduced in Chat-Scene [22] to associate each object with a unique learnable identifier token $<\!OBJ_{i}\!>$.
-These tokens are integrated into the tokenizer vocabulary and embedded alongside the embeddings of the object.
+To support unambiguous reference and interaction, we adopt the design introduced in Chat-Scene [huang2024chat] to associate each object with a unique learnable identifier token $OBJ_{i}\!>$.
+Concretely, each identifier token is inserted into the tokenizer vocabulary as a learnable lexical token and optimized end-to-end.
 3.3 Visual Embedding Extraction
 For each detected object, we extract geometric and visual embeddings from 3D point clouds and multi-view images.
-To capture 3D geometry, we use the pre-trained Uni3D [61] encoder, which provides high-quality spatial embeddings by processing each object’s segmented point cloud $\mathbf{P}_{i}$, producing an embedding vector $\mathbf{Z}^{v_{p}}_{i}\in\mathbb{R}^{1\times d}$.
-Uni3D is chosen for its strong performance in capturing fine-grained geometric structure across diverse 3D scenes.
-For 2D appearance, we employ the pre-trained DINOv2 [36] vision transformer to extract embeddings from multi-view images.
-The 3D mask of each object is projected onto these images, and the corresponding regions are cropped from the embedding maps and aggregated into a 2D visual embedding $\mathbf{Z}^{v_{f}}_{i}\in\mathbb{R}^{1\times d}$.
-DINOv2 is used for its ability to extract semantically rich embeddings from high-resolution visual input.
+For 3D geometry, we use the pre-trained Uni3D [zhou2023uni3d] encoder, which produces an embedding $\mathbf{Z}^{v_{p}}_{i}\in\mathbb{R}^{1\times d}$ from each segmented point cloud $\mathbf{P}_{i}$, and is known for capturing fine-grained structure across diverse scenes.
+For 2D appearance, we adopt the pre-trained DINOv2 [oquab2023dinov2] vision transformer to extract embeddings from multi-view images.
+We project each object’s 3D mask onto multi-view images and average the cropped DINOv2 embeddings per view.
+The features from different views are then aggregated using a size-weighted average into a 2D embedding $\mathbf{Z}^{v_{f}}_{i}\in\mathbb{R}^{1\times d}$.
+DINOv2 is chosen for its ability to extract semantically rich embeddings from high-resolution input.
 3.4 Text-Based Relational Modality
-While the embeddings extracted from both 2D multi-view images and 3D point clouds are powerful, these embeddings are inherently localized to individual objects.
-Specifically, 3D embeddings capture geometric and spatial attributes of each object in isolation, while 2D embeddings, though derived from full images, are ultimately aggregated per object and offer only limited relational context, making their relational awareness implicit and limited.
+While the embeddings extracted from both 2D multi-view images and 3D point clouds are powerful, they are inherently localized to individual objects.
+3D embeddings capture geometric and spatial attributes in isolation, while 2D embeddings, though derived from full images, are aggregated per object and provide only limited relational context, making their relational awareness implicit and limited.
 As a result, the visual representation lacks explicit modeling of inter-object interactions, limiting its capacity for comprehensive scene understanding and high-level reasoning.
 To overcome this limitation, we introduce a novel textual modality that encodes relational information in natural language.
-These descriptions explicitly capture how each object is situated within the scene and how it interacts with others, injecting rich contextual knowledge that is often inaccessible to vision-based embeddings alone.
-Our approach introduces this modality through a dual-level integration: (1) by encoding and fusing the text descriptions into the object representations and (2) by injecting them into the prompt space of the language model for enhanced reasoning.
-This design allows our model to take advantage of both low-level visual embeddings and high-level relational cues, ultimately improving performance on tasks that require contextual understanding.
+These descriptions capture how each object is situated within the scene and how it interacts with others, injecting contextual knowledge often inaccessible to vision-based embeddings alone.
+Our approach integrates this modality through a dual-level integration: (1) by encoding and fusing text descriptions into object representations and (2) by injecting them into the prompt space of the language model for enhanced reasoning.
+This design allows our model to exploit both low-level visual embeddings and high-level relational cues, ultimately improving performance on tasks requiring contextual understanding.
 Relational Description Generation.
@@
 Relational Descriptions Encoding.
-To encode the generated relational descriptions into the model, we use a Sentence-Transformer model [38] to convert each object’s description into a fixed-dimensional embedding vector $\mathbf{Z}^{v_{t}}_{i}$.
+To encode the generated relational descriptions into the model, we use a Sentence-Transformer model [reimers-2019-sentence-bert] to convert each object’s description into a fixed-dimensional embedding vector $\mathbf{Z}^{v_{t}}_{i}$.
 This representation is subsequently fused into the object representation, complementing the 2D visual embeddings $\mathbf{Z}^{v_{f}}_{i}$ and the 3D geometric embeddings $\mathbf{Z}^{v_{p}}_{i}$.
@@
 In addition to token-level fusion, we leverage a prompt-level strategy that delivers relational priors directly to the language model.
-When a query mentions a specific object by name or identifier, we prepend the description of that object to the input prompt.
+When a query mentions an object by name and contains a relational cue (e.g., “next to”), we append a brief description of the relevant object to the end of the input prompt.
 To avoid semantic ambiguity and ensure consistent grounding, we replace all object names within the descriptions with their corresponding object identifiers.
@@
 $\mathcal{L}(\theta)=-\sum_{i=1}^{k}\log P\left(s_{i}^{\text{res}}\mid s_{[1,\dots,i-1]}^{\text{res}},\,s^{\text{prefix}}\right),$ (3)
-where $k$ is the length of the response, $s_{i}^{\text{res}}$ is the generated target response, and $s_{[1,\dots,i-1]}^{\text{res}}$ is the previous $i-1$ tokens in the response.
+where $k$ is the length of the response, $s_{i}^{\text{res}}$ is the generated target response, and $s_{[1,\dots,i-1]}^{\text{res}}$ is the previous i1i-1 tokens in the response.
 $\theta$ denotes the trainable parameters.
 4 Experiments
-4.1 Datasets and Metrics
-Datasets.
-To evaluate the generalizability and task effectiveness of our approach, we conduct experiments on five benchmark datasets spanning four major 3D vision-language tasks.
-These include: ScanRefer [4] for single-object localization via visual grounding, Multi3DRefer [58] for compositional grounding of multiple targets, Scan2Cap [11] for 3D-aware caption generation, and both ScanQA [2] and SQA3D [34] for contextual question answering within 3D scenes.
-All five datasets are derived from the ScanNet dataset [13], a richly annotated collection of 1,513 indoor scenes featuring 3D point clouds, RGB images, and camera poses.
-This shared backbone ensures consistent scene semantics across benchmarks, facilitating multi-task learning.
-Each benchmark has been aligned to a unified format suitable for instruction-following scenarios.
-These datasets collectively evaluate a variety of 3D reasoning abilities, ranging from precise localization to compositional understanding.
-Metrics.
-We use standard evaluation protocols from prior work [22].
-For ScanRefer [4], we report thresholded accuracies Acc@0.25 and Acc@0.5, which assess whether the predicted object bounding box has an IoU with the ground truth exceeding 0.25 or 0.5.
-For Multi3DRefer [58], which involves grounding multiple targets, we use the F1 score at IoU thresholds of 0.25 and 0.5.
-For the captioning task Scan2Cap [11], we adopt CIDEr@0.5 and BLEU-4@0.5, integrating captioning quality with spatial alignment via IoU.
-For visual question answering, ScanQA [2] is evaluated using CIDEr and BLEU-4, while SQA3D [34] is evaluated using Exact Match (EM) and its refined variant EM-R as proposed in LEO [23].
-4.2 Implementation Details
-We extract 100 object proposals per scene using the Mask3D [39] segmentation model.
-For 3D geometry, we use Uni3D [61] to extract point embeddings $\mathbf{Z}^{v_{p}}_{i}$.
-For 2D appearance, we adopt DINOv2 [36] to extract 1024-dimensional embeddings per image.
-We project each object’s 3D mask onto multi-view images and average the cropped DINOv2 embeddings per view.
-These view-level embeddings are then aggregated using a size-weighted average to obtain the final object-level embeddings $\mathbf{Z}^{v_{f}}_{i}$.
-For text, we utilize the LLaVA-v1.5-7B [31] model to generate descriptions for each object.
-These object descriptions are then encoded using the all-mpnet-base-v2 model, a SentenceTransformer [38] based on MPNet [40] that captures rich contextual semantics for downstream fusion, resulting in the embeddings denoted as $\mathbf{Z}^{v_{t}}_{i}$.
-Each modality is projected to the token space using a three-layer MLP.
-We use Vicuna-7B-v1.5 [12] as the LLM and fine-tune it with LoRA [21] with a rank of 16.
-We train for 3 epochs using a batch size of 32, a base learning rate of $5\times 10^{-6}$, and a cosine annealing schedule.
-Training is conducted on two 80G NVIDIA A100 GPUs and completes in about 24 hours.
-We observe that training for 2 epochs yields improved performance on the ScanQA dataset, and we adopt this setting in relevant evaluations.
 Method ScanQA SQA3D ScanRefer
@@
 Expert Models
-ScanRefer [4] - - - - - - - 37.3 24.3
-ScanQA [2] 30.2 10.1 13.1 33.3 64.9 - - - -
-SQA3D [34] - - - - - 46.6 - - -
-3DJCG [3] - - - - - - - 49.6 37.3
-3D-VLP [27] 30.5 11.1 13.5 34.5 67.0 - - 51.4 39.5
-M3DRef-CLIP [58] - - - - - - - 51.9 44.7
-3D-VisTA [62] - 13.1 13.9 35.7 72.9 48.5 - 50.6 45.5
-ConcreteNet [41] - - - - - - - 50.6 46.5
-PQ3D [63] 43.0 - 17.8 - 87.8 47.1 - - 51.2
+ScanRefer [chen2020scanrefer] - - - - - - - 37.3 24.3
+ScanQA [azuma2022scanqa] 30.2 10.1 13.1 33.3 64.9 - - - -
+SQA3D [ma2022sqa3d] - - - - - 46.6 - - -
+3DJCG [cai20223djcg] - - - - - - - 49.6 37.3
+3D-VLP [jin2023context] 30.5 11.1 13.5 34.5 67.0 - - 51.4 39.5
+M3DRef-CLIP [zhang2023multi3drefer] - - - - - - - 51.9 44.7
+3D-VisTA [zhu20233d] - 13.1 13.9 35.7 72.9 48.5 - 50.6 45.5
+ConcreteNet [unal2023three] - - - - - - - 50.6 46.5
+PQ3D [zhu2024unifying] 43.0 - 17.8 - 87.8 47.1 - - 51.2
 LLM-based Models
-LAMM [52] - 5.8 - - 42.4 - - - 3.4
-Chat-3D [44] 29.1 6.4 11.9 28.5 53.2 - - - -
-ZSVG3D [54] - - - - - - - 36.4 32.7
-3D-LLM [20] 39.3 12.0 14.5 35.7 69.4 - - 30.3 -
-LL3DA [8] - 13.5 15.9 37.3 76.8 - - - -
-Grounded 3D-LLM [10] - 13.2 - - 75.9 - - 48.6 44.0
-LEO [23] - 11.5 16.2 39.3 80.0 50.0 52.4 - -
-Scene-LLM [14] 43.6 12.0 16.6 40.0 80.0 54.2 - - -
-3DGraphLLM [56] - 12.1 - - 87.6 53.1 - 57.0 51.3
-Chat-Scene [22] 43.2 14.3 18.0 41.6 87.7 54.6 57.5 55.5 50.2
-Descrip3D (Ours) 44.0 14.5 18.6 43.1 93.7 55.7 58.4 57.2 51.8
-Table 1: Performance on ScanQA [2], SQA3D [34], and ScanRefer [4].
+LAMM (Vicuna-13B) [yin2023lamm] - 5.8 - - 42.4 - - - 3.4
+Chat-3D (Vicuna-7B) [wang2023chat] 29.1 6.4 11.9 28.5 53.2 - - - -
+ZSVG3D (GPT-4) [yuan2024visual] - - - - - - - 36.4 32.7
+3D-LLM (BLIP-2-FlanT5) [hong20233d] 39.3 12.0 14.5 35.7 69.4 - - 30.3 -
+LL3DA (OPT-1.3B) [chen2024ll3da] - 13.5 15.9 37.3 76.8 - - - -
+Grounded 3D-LLM (Tiny-Vicuna-1B) [chen2024grounded] - 13.2 - - 75.9 - - 48.6 44.0
+LEO (Vicuna-7B) [huang2023embodied] - 11.5 16.2 39.3 80.0 50.0 52.4 - -
+Scene-LLM (LLaMA-2-7B) [fu2024scene] 43.6 12.0 16.6 40.0 80.0 54.2 - - -
+3DGraphLLM (Vicuna-7B) [zemskova20243dgraphllm] - 12.1 - - 87.6 53.1 - 57.0 51.3
+Chat-Scene (Vicuna-7B) [huang2024chat] 43.2 14.3 18.0 41.6 87.7 54.6 57.5 55.5 50.2
+Descrip3D (Ours, Vicuna-7B) 44.0 14.5 18.6 43.1 93.7 55.7 58.4 57.2 51.8
+Table 1: Performance on ScanQA [azuma2022scanqa], SQA3D [ma2022sqa3d], and ScanRefer [chen2020scanrefer].
 “Expert models” are tailored for specific tasks using task-oriented heads, while “LLM-based models” are designed for general instructions and responses.
 Descrip3D achieves the highest performance across all benchmarks, outperforming prior LLM-based methods by leveraging relational textual descriptions through a dual-level integration for more precise and context-aware reasoning.
+(a) Qualitative Comparison of Question Answering.
+(b) Qualitative Comparison of Grounding.
+Figure 3: Qualitative comparison of 3D scene understanding tasks.
+Descrip3D outperforms Chat-Scene, especially in cases involving complex spatial grounding or multi-object reasoning, due to its use of a dual-level integrated relational textual descriptions that enhance contextual understanding.
+4.1 Datasets and Metrics
+Datasets.
+To evaluate generalizability, we conduct experiments on five benchmarks spanning four major 3D vision-language tasks: ScanRefer [chen2020scanrefer] for single-object grounding, Multi3DRefer [zhang2023multi3drefer] for compositional multi-object grounding, Scan2Cap [chen2021scan2cap] for 3D-aware captioning, and ScanQA [azuma2022scanqa] plus SQA3D [ma2022sqa3d] for scene-level QA.
+All are derived from ScanNet [dai2017scannet], a richly annotated dataset of 1,513 indoor scenes with point clouds, RGB images, and camera poses.
+This shared backbone provides consistent semantics across tasks and enables multi-task evaluation.
+In our experiments, we strictly follow the official splits provided by each benchmark (ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D).
+All are built on ScanNet, where splits are defined at the scene level, ensuring disjoint train/val/test sets.
+Thus, evaluation scenes and objects are never partially seen during training, eliminating data leakage.
+Metrics.
+We use standard evaluation protocols from prior work [huang2024chat].
+For ScanRefer [chen2020scanrefer], we report thresholded accuracies Acc@0.25 and Acc@0.5, which assess whether the predicted object bounding box has an IoU with the ground truth exceeding 0.25 or 0.5.
+For Multi3DRefer [zhang2023multi3drefer], which involves grounding multiple targets, we use the F1 score at IoU thresholds of 0.25 and 0.5.
+For the captioning task Scan2Cap [chen2021scan2cap], we adopt CIDEr@0.5 and BLEU-4@0.5, integrating captioning quality with spatial alignment via IoU.
+For visual question answering, ScanQA [azuma2022scanqa] is evaluated using CIDEr and BLEU-4, while SQA3D [ma2022sqa3d] is evaluated using Exact Match (EM) and its refined variant EM-R as proposed in LEO [huang2023embodied].
+3D 2D Text ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
+Acc@0.5 F1@0.5 C@0.5 CIDEr EM
+42.0 44.3 63.5 86.5 52.7
+44.0 48.1 68.4 86.2 52.4
+2.0 8.2 16.1 73.2 49.7
+46.7 49.5 73.6 89.4 53.8
+41.5 44.9 63.3 86.9 53.1
+49.5 52.8 73.0 90.4 54.1
+51.8 55.1 77.2 93.7 55.7
+Table 2: Modality ablation results across five benchmarks.
+Text-only features underperform due to lack of geometric grounding, while 2D and 3D features alone offer moderate performance.
+The best results are achieved when all three modalities are integrated, highlighting their complementary strengths.
+Method Multi3DRefer
+F1@0.25 F1@0.5
+Expert Models
+3DVG-Transformer [zhao20213dvg] - 25.5
+3DJCG [cai20223djcg] - 26.6
+D3Net [chen2022d3net] - 32.2
+M3DRef-CLIP [zhang2023multi3drefer] 42.8 38.4
+PQ3D [zhu2024unifying] - 50.1
+LLM-based Models
+Grounded 3D-LLM [chen2024grounded] 44.7 40.8
+Chat-Scene [huang2024chat] 57.1 52.4
+Descrip3D (Ours) 59.4 55.1
+Table 3: Performance on Multi3DRefer [zhang2023multi3drefer]. Descrip3D achieves the best performance, with notable gains in multi-object reasoning due to its dual-level relational text modeling.
+Method Scan2Cap
+C@0.5 B-4@0.5
+Expert Models
+Scan2Cap [chen2021scan2cap] 35.2 22.4
+3DJCG [cai20223djcg] 49.5 31.0
+3D-VLP [jin2023context] 54.9 32.3
+3D-VisTA [zhu20233d] 66.9 34.0
+Vote2Cap-DETR++ [chen2024vote2cap] 67.6 37.1
+LLM-based Models
+LL3DA [chen2024ll3da] 65.2 36.8
+LEO [huang2023embodied] 68.4 36.9
+Grounded 3D-LLM [chen2024grounded] 70.2 35.0
+Chat-Scene [huang2024chat] 77.1 36.3
+Descrip3D (Ours) 77.2 34.5
+Table 4: Performance on Scan2Cap [chen2021scan2cap]. Descrip3D achieves the best C@0.5 performance by leveraging richer object-level relational context.
+4.2 Implementation Details
+We extract 100 object proposals per scene using the Mask3D [schult2022mask3d] segmentation model.
+For 3D geometry, we use Uni3D [zhou2023uni3d] to extract point embeddings $\mathbf{Z}^{v_{p}}_{i}$.
+For 2D appearance, we adopt DINOv2 [oquab2023dinov2] to extract object-level embeddings $\mathbf{Z}^{v_{f}}_{i}$, consistent with prior work (e.g., ChatScene [huang2024chat], 3DGraphLLM [zemskova20243dgraphllm]), to ensure fair comparison.
+While stronger encoders could raise performance, our relational-text modality is independent of the 2D backbone.
+For text, we utilize the LLaVA-v1.5-7B [liu2023llava] model to generate descriptions for each object.
+These object descriptions are then encoded using the all-mpnet-base-v2 model, a SentenceTransformer [reimers-2019-sentence-bert] based on MPNet [song2020mpnet] that captures rich contextual semantics for downstream fusion, resulting in the embeddings denoted as $\mathbf{Z}^{v_{t}}_{i}$.
+Each modality is projected to the token space using a three-layer MLP.
+We use Vicuna-7B-v1.5 [chiang2023vicuna] as the base LLM to ensure a fair comparison with prior work (ChatScene, 3DGraphLLM, 3D-Ground-LLM).
+Grounded-3D-LLM primarily reports results with Tiny-Vicuna-1B, but also notes little difference with Vicuna-7B.
+We fine-tune it with LoRA [hu2022lora] (rank 16) for 3 epochs, batch size 32, lr $5\times 10^{-6}$, cosine schedule.
+Training is conducted on two 80G NVIDIA A100 GPUs and completes in about 24 hours.
+We observe that training for 2 epochs yields improved performance on the ScanQA dataset, and we adopt this setting in relevant evaluations.
 4.3 Performance Comparison
@@
 These tasks are benchmarked using the five widely adopted datasets mentioned before.
+Reference Style ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
+Acc@0.5 F1@0.5 C@0.5 CIDEr EM
+Object Name Only 51.6 54.5 75.6 91.4 55.1
+Object Name + ID 51.1 53.9 74.1 91.7 54.2
+Object ID Only (Ours) 51.8 55.1 77.2 93.7 55.7
+Table 5: Ablation study on object reference style in prompt-level text injection.
+Using object IDs alone yields the best performance across all benchmarks, as it avoids ambiguity and improves alignment between queries and object descriptions.
+Embedding Prompt ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
+Acc@0.5 F1@0.5 C@0.5 CIDEr EM
+50.2 52.4 77.1 87.7 54.6
+50.9 54.7 76.4 93.6 55.2
+51.6 55.1 77.3 89.9 54.7
+51.8 55.1 77.2 93.7 55.7
+Table 6: Ablation study on text description and injection strategy.
+Combining embedding-level and prompt-level injection consistently leads to the best overall performance, demonstrating the complementary benefits of dual-level relational text integration.
 Scene-level Question Answering.
-We evaluate our model on ScanQA [2] and SQA3D [34], both of which require comprehensive scene-level understanding and precise object-grounded reasoning.
-As shown in Table 1, our method achieves the highest scores across all major metrics compared with previous expert models and LLM-based methods, highlighting the strength of our dual-level integrated relational descriptions in enhancing context understanding.
+We evaluate our model on ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d], both of which require comprehensive scene-level understanding and precise object-grounded reasoning.
+As shown in table˜1, our method achieves the highest scores across all major metrics compared with previous expert models and LLM-based methods, highlighting the strength of our dual-level integrated relational descriptions in enhancing context understanding.
 Notably, CIDEr, which emphasizes content relevance, shows a significant gain of 5.9 over PQ3D and 6.0 over Chat-Scene, confirming that our model generates more informative and relevant answers.
@@
 3D Visual Grounding.
-We evaluate 3D visual grounding on the ScanRefer [4] dataset.
-As shown in Table 1, our method achieves the best performance among all methods.
+We evaluate 3D visual grounding on the ScanRefer [chen2020scanrefer] dataset.
+As shown in table˜1, our method achieves the best performance among all methods.
 Compared to strong expert models like ConcreteNet and 3D-VisTA, our model surpasses them by 5.3 and 6.3 at the stricter 0.5 IoU threshold, respectively.
@@
 The improvement over Chat-Scene, which uses only 2D and 3D embeddings, validates the great benefit of incorporating relational object-level textual descriptions through a dual-level integration approach.
-Multi-Object 3D Visual Grounding.
-We evaluate our method on Multi3DRefer [58], a challenging benchmark requiring models to resolve complex multi-object visual grounding within 3D scenes.
+Multi‑Object 3D Visual Grounding.
+We evaluate our method on Multi3DRefer [zhang2023multi3drefer], a challenging benchmark requiring models to resolve complex multi-object visual grounding within 3D scenes.
 Unlike traditional single-object grounding, this task demands fine-grained relational reasoning among multiple entities.
-As shown in Table 2, our method achieves the best performance, outperforming both expert models and LLM-based baselines such as Chat-Scene and Grounded 3D-LLM.
-Method Multi3DRefer
-F1@0.25 F1@0.5
-Expert Models
-3DVG-Transformer [59] - 25.5
-3DJCG [3] - 26.6
-D3Net [5] - 32.2
-M3DRef-CLIP [58] 42.8 38.4
-PQ3D [63] - 50.1
-LLM-based Models
-Grounded 3D-LLM [10] 44.7 40.8
-Chat-Scene [22] 57.1 52.4
-Descrip3D (Ours) 59.4 55.1
-Table 2: Performance on Multi3DRefer [58]. Descrip3D achieves the best performance, with notable gains in multi-object reasoning due to its dual-level relational text modeling.
+As shown in table˜3, our method achieves the best performance, outperforming both expert models and LLM-based baselines such as Chat-Scene and Grounded 3D-LLM.
 3D Captioning.
-We evaluate our method on the Scan2Cap [11] benchmark, which focuses on generating natural language descriptions for 3D objects in complex indoor scenes.
-As shown in Table 3, our method achieves the highest CIDEr score, indicating that the generated captions are more informative and semantically aligned with the ground truth.
-While our BLEU-4 score is lower than some other models, this metric emphasizes exact n-gram overlap and can penalize variation in phrasing, particularly in models like ours that incorporate diverse relational language.
-Importantly, Descrip3D is trained as a unified model across multiple tasks without task-specific tuning, unlike expert models tailored for captioning.
-Despite this, it still delivers a strong overall performance, suggesting that our relational text descriptions effectively enhance semantic grounding and generalization across diverse scene-language tasks.
-Method Scan2Cap
-C@0.5 B-4@0.5
-Expert Models
-Scan2Cap [11] 35.2 22.4
-3DJCG [3] 49.5 31.0
-3D-VLP [27] 54.9 32.3
-3D-VisTA [62] 66.9 34.0
-Vote2Cap-DETR++ [9] 67.6 37.1
-LLM-based Models
-LL3DA [8] 65.2 36.8
-LEO [23] 68.4 36.9
-Grounded 3D-LLM [10] 70.2 35.0
-Chat-Scene [22] 77.1 36.3
-Descrip3D (Ours) 77.2 34.5
-Table 3: Performance on Scan2Cap [11]. Descrip3D achieves the best C@0.5 performance by leveraging richer object-level relational context.
+On Scan2Cap [chen2021scan2cap], which evaluates the generation of 3D-aware captions for objects in complex indoor scenes, our model obtains the highest CIDEr score as shown in table˜4.
+This indicates that its captions are more informative and semantically aligned with ground truth.
+While BLEU-4 is slightly lower, this metric favors exact n-gram overlap and can penalize the diverse relational phrasing produced by our model.
+Importantly, unlike expert captioning systems tailored to this task, Descrip3D is trained as a unified model across multiple benchmarks yet still delivers strong captioning results.
+This shows that relational textual descriptions not only improve grounding and QA but also enhance semantic richness and generalization.
 Qualitative results.
-Figure 3 presents qualitative comparisons between our method and Chat-Scene on two representative tasks: 3D question answering (Figure 3(a)) and 3D visual grounding (Figure 3(b)).
-For question answering, our model demonstrates a stronger ability to understand relational context and resolve spatial references.
-In comparison, Chat-Scene demonstrably struggles more with object relationships.
-For visual grounding, our model more accurately identifies target objects based on dual-level integrated complex natural language descriptions that involve appearance and relative position.
-In contrast, Chat-Scene frequently fails to disambiguate between visually or semantically similar candidates.
+fig.˜3 presents qualitative comparisons between our method and Chat-Scene on two representative tasks: 3D question answering (fig.˜3(a)) and 3D visual grounding (fig.˜3(b)).
+For question answering, our model shows stronger ability to understand relational context and resolve spatial references, whereas Chat-Scene struggles more with object relationships.
+For visual grounding, our model more accurately identifies target objects based on dual-level integrated complex natural language descriptions involving appearance and relative position, while Chat-Scene often fails to disambiguate visually or semantically similar candidates.
 These results highlight the benefit of injecting detailed object-level descriptions into the language model through a dual-level approach, enabling more precise and interpretable multi-modal reasoning.
-(a) Qualitative Comparison of Question Answering.
-(b) Qualitative Comparison of Grounding.
-Figure 3: Qualitative comparison of 3D scene understanding tasks.
-Descrip3D outperforms Chat-Scene, especially in cases involving complex spatial grounding or multi-object reasoning, due to its use of a dual-level integrated relational textual descriptions that enhance contextual understanding.
-In summary, our method achieves strong and consistent results across the five benchmarks, demonstrating its effectiveness in different tasks.
+In summary, our method achieves strong and consistent results across five benchmarks, demonstrating its effectiveness across tasks.
 4.4 Ablation Studies
-To evaluate the impact of our proposed object-level text descriptions and the dual-level integration strategies, we conduct a series of ablation experiments across key components of our model.
+To evaluate the impact of our object-level text descriptions and dual-level integration strategies, we perform ablations on key components of the model.
+Effect of Individual Modalities.
+We ablate 3D, 2D, and text features as shown in table˜2.
+Single modalities are limited: 3D captures geometry but lacks appearance and relations, 2D provides surface cues without spatial grounding, and text alone conveys semantics but cannot localize objects.
+Pairwise combinations yield complementary gains, while integrating all three produces the strongest results.
+This confirms that relational text is most effective when fused with both 2D and 3D features.
 Effect of Object Reference Style in Prompt-Level Description.
-We examine how the choice of object reference expression in prompt-level text injection affects performance.
-As shown in Table 4, using raw object names can lead to suboptimal performance due to ambiguity when multiple instances of the same category are present.
-Appending object IDs to names may confuse the model further, failing to improve performance and potentially introducing redundancy.
-In contrast, using object IDs alone yields the best results across all benchmarks.
-This strategy eliminates ambiguity and ensures precise alignment between the referenced objects in the query and their corresponding descriptions.
-Reference Style ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
-Acc@0.5 F1@0.5 C@0.5 CIDEr EM
-Object Name Only 51.6 54.5 75.6 91.4 55.1
-Object Name + ID 51.1 53.9 74.1 91.7 54.2
-Object ID Only (Ours) 51.8 55.1 77.2 93.7 55.7
-Table 4: Ablation study on object reference style in prompt-level text injection.
-Using object IDs alone yields the best performance across all benchmarks, as it avoids ambiguity and improves alignment between queries and object descriptions.
+We test different ways of expressing objects in prompt-level text injection as shown in table˜5.
+Using raw object names leads to ambiguity when multiple instances share the same category.
+Adding IDs to names introduces redundancy without improvement.
+In contrast, using IDs alone consistently gives the best results, ensuring unambiguous alignment between queries and object descriptions across all benchmarks.
 Effect of Object Descriptions and Fusion Location.
-We study the impact of incorporating object-level textual descriptions at different points in the model.
-As shown in Table 5, injecting textual descriptions solely as embeddings improves grounding and question answering by enriching object semantics.
-Injecting them into the prompt also boosts performance on tasks like ScanQA and Scan2Cap.
-Combining both yields the best results on most datasets, highlighting their complementarity.
-However, on Multi3DRefer, this strategy slightly underperforms, possibly due to redundant signals in complex scenes.
-Overall, the dual-level integration enhances both spatial grounding and semantic reasoning.
-Embedding Prompt ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
-Acc@0.5 F1@0.5 C@0.5 CIDEr EM
-50.2 52.4 77.1 87.7 54.6
-51.5 55.4 74.6 89.3 54.5
-51.7 55.5 75.3 92.3 53.6
-51.8 55.1 77.2 93.7 55.7
-Table 5: Ablation study on text description and injection strategy.
-Combining embedding-level and prompt-level injection consistently leads to the best overall performance, demonstrating the complementary benefits of dual-level relational text integration.
-5 Conclusion and Future Work
+We also study where textual descriptions are integrated, as shown in table˜6.
+Injecting them only as embeddings enriches object semantics and improves QA tasks.
+Injecting them only into prompts benefits tasks like Scan2Cap and Multi3DRefer.
+Combining both yields the best overall results across all benchmarks, with especially strong improvements on ScanQA.
+These findings show that embedding- and prompt-level strategies are complementary, and together provide the most robust improvements in both spatial grounding and semantic reasoning.
+5 Conclusion
 We propose Descrip3D, a simple yet powerful framework for 3D scene understanding that explicitly models inter-object relationships using object-level textual descriptions.
@@
 As future work, we plan to explore end-to-end training that jointly learns object description generation and extend our approach to dynamic or outdoor environments.
-Reference
-  • [1] Panos Achlioptas, Ahmed Abdelreheem, Fei Xia, Mohamed Elhoseiny, and Leonidas Guibas.
-Referit3d: Neural listeners for fine-grained 3d object identification in real-world scenes.
-In European conference on computer vision, pages 422–440.
-Springer, 2020.
-  • [2] Daichi Azuma, Taiki Miyanishi, Shuhei Kurita, and Motoaki Kawanabe.
-Scanqa: 3d question answering for spatial scene understanding.
-In proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19129–19139, 2022.
-  • [3] Daigang Cai, Lichen Zhao, Jing Zhang, Lu Sheng, and Dong Xu.
-3djcg: A unified framework for joint dense captioning and visual grounding on 3d point clouds.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 16464–16473, 2022.
-  • [4] Dave Zhenyu Chen, Angel X Chang, and Matthias Nießner.
-Scanrefer: 3d object localization in rgb-d scans using natural language.
-In European conference on computer vision, pages 202–221.
-Springer, 2020.
-  • [5] Dave Zhenyu Chen, Qirui Wu, Matthias Nießner, and Angel X. Chang. D3net: A speaker-listener architecture for semi-supervised dense captioning and visual grounding in rgb-d scans, 2021.
-  • [6] Jiaming Chen, Weixin Luo, Xiaolin Wei, Lin Ma, and Wei Zhang. Ham: Hierarchical attention model with high performance for 3d visual grounding. arXiv preprint arXiv:2210.12513, 2(3), 2022.
-  • [7] Shizhe Chen, Pierre-Louis Guhur, Makarand Tapaswi, Cordelia Schmid, and Ivan Laptev.
-Language conditioned spatial relation reasoning for 3d object grounding.
-Advances in neural information processing systems, 35:20522–20535, 2022.
-  • [8] Sijin Chen, Xin Chen, Chi Zhang, Mingsheng Li, Gang Yu, Hao Fei, Hongyuan Zhu, Jiayuan Fan, and Tao Chen.
-Ll3da: Visual interactive instruction tuning for omni-3d understanding reasoning and planning.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 26428–26438, 2024.
-  • [9] Sijin Chen, Hongyuan Zhu, Mingsheng Li, Xin Chen, Peng Guo, Yinjie Lei, Gang Yu, Taihao Li, and Tao Chen.
-Vote2cap-detr++: Decoupling localization and describing for end-to-end 3d dense captioning.
-IEEE Transactions on Pattern Analysis and Machine Intelligence, 46(11):7331–7347, 2024.
-  • [10] Yilun Chen, Shuai Yang, Haifeng Huang, Tai Wang, Runsen Xu, Ruiyuan Lyu, Dahua Lin, and Jiangmiao Pang. Grounded 3d-llm with referent tokens. arXiv preprint arXiv:2405.10370, 2024.
-  • [11] Zhenyu Chen, Ali Gholami, Matthias Nießner, and Angel X Chang.
-Scan2cap: Context-aware dense captioning in rgb-d scans.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 3193–3203, 2021.
-  • [12] Wei-Lin Chiang, Zhuohan Li, Ziqing Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E Gonzalez, et al.
-Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality.
-See https://vicuna.
-lmsys.
-org (accessed 14 April 2023), 2(3):6, 2023.
-  • [13] Angela Dai, Angel X Chang, Manolis Savva, Maciej Halber, Thomas Funkhouser, and Matthias Nießner.
-Scannet: Richly-annotated 3d reconstructions of indoor scenes.
-In Proceedings of the IEEE conference on computer vision and pattern recognition, pages 5828–5839, 2017.
-  • [14] Rao Fu, Jingyu Liu, Xilun Chen, Yixin Nie, and Wenhan Xiong. Scene-llm: Extending language model for 3d visual understanding and reasoning. arXiv preprint arXiv:2403.11401, 2024.
-  • [15] Qiao Gu, Ali Kuwajerwala, Sacha Morin, Krishna Murthy Jatavallabhula, Bipasha Sen, Aditya Agarwal, Corban Rivera, William Paul, Kirsty Ellis, Rama Chellappa, et al.
-Conceptgraphs: Open-vocabulary 3d scene graphs for perception and planning.
-In 2024 IEEE International Conference on Robotics and Automation (ICRA), pages 5021–5028.
-IEEE, 2024.
-  • [16] Ziyu Guo, Renrui Zhang, Xiangyang Zhu, Yiwen Tang, Xianzheng Ma, Jiaming Han, Kexin Chen, Peng Gao, Xianzhi Li, Hongsheng Li, et al.
-Point-bind & point-llm: Aligning point cloud with multi-modality for 3d understanding, generation, and instruction following.
-arXiv preprint arXiv:2309.00615, 2023.
-  • [17] Muhammad Usman Hadi, Rizwan Qureshi, Abbas Shah, Muhammad Irfan, Anas Zafar, Muhammad Bilal Shaikh, Naveed Akhtar, Jia Wu, Seyedali Mirjalili, et al.
-A survey on large language models: Applications, challenges, limitations, and practical usage.
-Authorea Preprints, 2023.
-  • [18] Jiaming Han, Renrui Zhang, Wenqi Shao, Peng Gao, Peng Xu, Han Xiao, Kaipeng Zhang, Chris Liu, Song Wen, Ziyu Guo, et al.
-Imagebind-llm: Multi-modality instruction tuning.
-arXiv preprint arXiv:2309.03905, 2023.
-  • [19] Deepti Hegde, Jeya Maria Jose Valanarasu, and Vishal Patel.
-Clip goes 3d: Leveraging prompt tuning for language grounded 3d recognition.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2028–2038, 2023.
-  • [20] Yining Hong, Haoyu Zhen, Peihao Chen, Shuhong Zheng, Yilun Du, Zhenfang Chen, and Chuang Gan.
-3d-llm: Injecting the 3d world into large language models.
-Advances in Neural Information Processing Systems, 36:20482–20494, 2023.
-  • [21] Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, Weizhu Chen, et al. Lora: Low-rank adaptation of large language models. ICLR, 1(2):3, 2022.
-  • [22] Haifeng Huang, Yilun Chen, Zehan Wang, Rongjie Huang, Runsen Xu, Tai Wang, Luping Liu, Xize Cheng, Yang Zhao, Jiangmiao Pang, et al.
-Chat-scene: Bridging 3d scene and large language models with object identifiers.
-Advances in Neural Information Processing Systems, 37:113991–114017, 2024.
-  • [23] Jiangyong Huang, Silong Yong, Xiaojian Ma, Xiongkun Linghu, Puhao Li, Yan Wang, Qing Li, Song-Chun Zhu, Baoxiong Jia, and Siyuan Huang.
-An embodied generalist agent in 3d world.
-arXiv preprint arXiv:2311.12871, 2023.
-  • [24] Pin-Hao Huang, Han-Hung Lee, Hwann-Tzong Chen, and Tyng-Luh Liu.
-Text-guided graph neural networks for referring 3d instance segmentation.
-In Proceedings of the AAAI Conference on Artificial Intelligence, volume 35, pages 1610–1618, 2021.
-  • [25] Shijia Huang, Yilun Chen, Jiaya Jia, and Liwei Wang.
-Multi-view transformer for 3d visual grounding.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 15524–15533, 2022.
-  • [26] Ayush Jain, Nikolaos Gkanatsios, Ishita Mediratta, and Katerina Fragkiadaki.
-Bottom up top down detection transformers for language grounding in images and point clouds.
-In European Conference on Computer Vision, pages 417–433.
-Springer, 2022.
-  • [27] Zhao Jin, Munawar Hayat, Yuwei Yang, Yulan Guo, and Yinjie Lei.
-Context-aware alignment and mutual masking for 3d-language pre-training.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 10984–10994, 2023.
-  • [28] Xinyi Li, Sai Wang, Siqi Zeng, Yu Wu, and Yi Yang. A survey on llm-based multi-agent systems: workflow, infrastructure, and challenges. Vicinagearth, 1(1):9, 2024.
-  • [29] Zechuan Li, Hongshan Yu, Yihao Ding, Yan Li, Yong He, and Naveed Akhtar.
-Embodied intelligence for 3d understanding: A survey on 3d scene question answering.
-arXiv preprint arXiv:2502.00342, 2025.
-  • [30] Daizong Liu, Yang Liu, Wencan Huang, and Wei Hu. A survey on text-guided 3d visual grounding: Elements, recent advances, and future directions. arXiv preprint arXiv:2406.05785, 2024.
-  • [31] Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. Visual instruction tuning. In NeurIPS, 2023.
-  • [32] Junyu Luo, Jiahui Fu, Xianghao Kong, Chen Gao, Haibing Ren, Hao Shen, Huaxia Xia, and Si Liu.
-3d-sps: Single-stage 3d visual grounding via referred point progressive selection.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 16454–16463, 2022.
-  • [33] Xianzheng Ma, Yash Bhalgat, Brandon Smart, Shuai Chen, Xinghui Li, Jian Ding, Jindong Gu, Dave Zhenyu Chen, Songyou Peng, Jia-Wang Bian, et al.
-When llms step into the 3d world: A survey and meta-analysis of 3d tasks via multi-modal large language models.
-arXiv preprint arXiv:2405.10255, 2024.
-  • [34] Xiaojian Ma, Silong Yong, Zilong Zheng, Qing Li, Yitao Liang, Song-Chun Zhu, and Siyuan Huang. Sqa3d: Situated question answering in 3d scenes. arXiv preprint arXiv:2210.07474, 2022.
-  • [35] Shervin Minaee, Tomas Mikolov, Narjes Nikzad, Meysam Chenaghlu, Richard Socher, Xavier Amatriain, and Jianfeng Gao. Large language models: A survey. arXiv preprint arXiv:2402.06196, 2024.
-  • [36] Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al.
-Dinov2: Learning robust visual features without supervision.
-arXiv preprint arXiv:2304.07193, 2023.
-  • [37] Zhangyang Qi, Zhixiong Zhang, Ye Fang, Jiaqi Wang, and Hengshuang Zhao. Gpt4scene: Understand 3d scenes from videos with vision-language models. arXiv preprint arXiv:2501.01428, 2025.
-  • [38] Nils Reimers and Iryna Gurevych.
-Sentence-bert: Sentence embeddings using siamese bert-networks.
-In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing.
-Association for Computational Linguistics, 11 2019.
-  • [39] Jonas Schult, Francis Engelmann, Alexander Hermans, Or Litany, Siyu Tang, and Bastian Leibe.
-Mask3d: Mask transformer for 3d semantic instance segmentation.
-arXiv preprint arXiv:2210.03105, 2022.
-  • [40] Kaitao Song, Xu Tan, Tao Qin, Jianfeng Lu, and Tie-Yan Liu.
-Mpnet: Masked and permuted pre-training for language understanding.
-Advances in neural information processing systems, 33:16857–16867, 2020.
-  • [41] Ozan Unal, Christos Sakaridis, Suman Saha, Fisher Yu, and Luc Van Gool. Three ways to improve verbo-visual fusion for dense 3d visual grounding. arXiv preprint arXiv:2309.04561, 2:15, 2023.
-  • [42] Yun-Cheng Wang, Jintang Xue, Chengwei Wei, and C-C Jay Kuo.
-An overview on generative ai at scale with edge–cloud computing.
-IEEE Open Journal of the Communications Society, 4:2952–2971, 2023.
-  • [43] Zehan Wang, Haifeng Huang, Yang Zhao, Linjun Li, Xize Cheng, Yichen Zhu, Aoxiong Yin, and Zhou Zhao.
-3drp-net: 3d relative position-aware network for 3d visual grounding.
-arXiv preprint arXiv:2307.13363, 2023.
-  • [44] Zehan Wang, Haifeng Huang, Yang Zhao, Ziang Zhang, and Zhou Zhao.
-Chat-3d: Data-efficiently tuning large language model for universal dialogue of 3d scenes.
-arXiv preprint arXiv:2308.08769, 2023.
-  • [45] Abdelrhman Werby, Chenguang Huang, Martin Büchner, Abhinav Valada, and Wolfram Burgard.
-Hierarchical open-vocabulary 3d scene graphs for language-grounded robot navigation.
-In First Workshop on Vision-Language Models for Navigation and Manipulation at ICRA 2024, 2024.
-  • [46] Tung-Yu Wu, Sheng-Yu Huang, and Yu-Chiang Frank Wang. Dora: 3d visual grounding with order-aware referring. CoRR, 2024.
-  • [47] Yanmin Wu, Xinhua Cheng, Renrui Zhang, Zesen Cheng, and Jian Zhang.
-Eda: Explicit text-decoupling and dense alignment for 3d visual grounding.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19231–19242, 2023.
-  • [48] Runsen Xu, Xiaolong Wang, Tai Wang, Yilun Chen, Jiangmiao Pang, and Dahua Lin.
-Pointllm: Empowering large language models to understand point clouds.
-In European Conference on Computer Vision, pages 131–147.
-Springer, 2024.
-  • [49] Jintang Xue, Yun-Cheng Wang, Chengwei Wei, Xiaofeng Liu, Jonghye Woo, C-C Jay Kuo, et al.
-Bias and fairness in chatbots: An overview.
-APSIPA Transactions on Signal and Information Processing, 13(2), 2024.
-  • [50] Le Xue, Mingfei Gao, Chen Xing, Roberto Martín-Martín, Jiajun Wu, Caiming Xiong, Ran Xu, Juan Carlos Niebles, and Silvio Savarese.
-Ulip: Learning a unified representation of language, images, and point clouds for 3d understanding.
-In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 1179–1189, 2023.
-  • [51] Zhengyuan Yang, Songyang Zhang, Liwei Wang, and Jiebo Luo.
-Sat: 2d semantics assisted training for 3d visual grounding.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1856–1866, 2021.
-  • [52] Zhenfei Yin, Jiong Wang, Jianjian Cao, Zhelun Shi, Dingning Liu, Mukai Li, Xiaoshui Huang, Zhiyong Wang, Lu Sheng, Lei Bai, et al.
-Lamm: Language-assisted multi-modal instruction-tuning dataset, framework, and benchmark.
-Advances in Neural Information Processing Systems, 36:26650–26685, 2023.
-  • [53] Ting Yu, Xiaojun Lin, Shuhui Wang, Weiguo Sheng, Qingming Huang, and Jun Yu.
-A comprehensive survey of 3d dense captioning: Localizing and describing objects in 3d scenes.
-IEEE Transactions on Circuits and Systems for Video Technology, 34(3):1322–1338, 2023.
-  • [54] Zhihao Yuan, Jinke Ren, Chun-Mei Feng, Hengshuang Zhao, Shuguang Cui, and Zhen Li.
-Visual programming for zero-shot open-vocabulary 3d visual grounding.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 20623–20633, 2024.
-  • [55] Zhihao Yuan, Xu Yan, Yinghong Liao, Ruimao Zhang, Sheng Wang, Zhen Li, and Shuguang Cui.
-Instancerefer: Cooperative holistic understanding for visual grounding on point clouds through instance multi-level contextual referring.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1791–1800, 2021.
-  • [56] Tatiana Zemskova and Dmitry Yudin. 3dgraphllm: Combining semantic graphs and large language models for 3d scene understanding. arXiv preprint arXiv:2412.18450, 2024.
-  • [57] Jirong Zha, Yuxuan Fan, Xiao Yang, Chen Gao, and Xinlei Chen. How to enable llm with 3d capacity? a survey of spatial reasoning in llm. arXiv preprint arXiv:2504.05786, 2025.
-  • [58] Yiming Zhang, ZeMing Gong, and Angel X Chang.
-Multi3drefer: Grounding text description to multiple 3d objects.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 15225–15236, 2023.
-  • [59] Lichen Zhao, Daigang Cai, Lu Sheng, and Dong Xu.
-3dvg-transformer: Relation modeling for visual grounding on point clouds.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2928–2937, 2021.
-  • [60] Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al.
-A survey of large language models.
-arXiv preprint arXiv:2303.18223, 1(2), 2023.
-  • [61] Junsheng Zhou, Jinsheng Wang, Baorui Ma, Yu-Shen Liu, Tiejun Huang, and Xinlong Wang. Uni3d: Exploring unified 3d representation at scale. arXiv preprint arXiv:2310.06773, 2023.
-  • [62] Ziyu Zhu, Xiaojian Ma, Yixin Chen, Zhidong Deng, Siyuan Huang, and Qing Li.
-3d-vista: Pre-trained transformer for 3d vision and text alignment.
-In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2911–2921, 2023.
-  • [63] Ziyu Zhu, Zhuofan Zhang, Xiaojian Ma, Xuesong Niu, Yixin Chen, Baoxiong Jia, Zhidong Deng, Siyuan Huang, and Qing Li.
-Unifying 3d vision-language understanding via promptable queries.
-In European Conference on Computer Vision, pages 188–206.
-Springer, 2024.
+6 Acknowledgments
+This project was sponsored by the US DoD LUCI (Laboratory University Collaboration Initiative) fellowship and the US Army Research Laboratory.
+The authors also acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing computing resources.
 Appendix A Examples of Generated Descriptions
-Figure 4: Qualitative Examples of Object-level Relational Descriptions Generated from Multi-view Images.
-The upper left part displays the image with object names, and the lower left part shows the bounding boxes of the objects.
-We present qualitative examples of our object-level textual descriptions in Figure 4.
+Figure 4: Qualitative examples of object-level relational descriptions generated using Prompt A (Default) with LLaVA-1.5
+.
+The upper left part displays the image with object names, and the lower left shows the 3D bounding boxes.
+The right side contains relational descriptions emphasizing general proximity and object co-occurrence without enforcing strict spatial language.
+Figure 5: Qualitative examples of object-level relational descriptions generated using Prompt B (Spatially Focused) with LLaVA-1.5.
+Compared to Prompt A, these descriptions include more explicit spatial terms (e.g., “on the left,” “behind”) and visual attributes, resulting in shorter but more positionally grounded sentences.
+We present qualitative examples of our object-level textual descriptions in  fig.˜4.
 Starting from detected object proposals and their corresponding multi-view images, we overlay the object names at the center of the projection areas in the image, as illustrated in the upper left part of the example.
@@
 To examine the impact of explicitly overlaying object category names during relational description generation, we conduct an ablation study comparing two variants: one where multi-view images include projected object labels (ours), and one without.
-As shown in Table 6, incorporating object labels consistently improves performance across all five benchmarks.
+As shown in table˜7, incorporating object labels consistently improves performance across all five benchmarks.
 The improvement is particularly notable in Scan2Cap and SQA3D, where more precise object references in the descriptions likely benefit caption generation and question answering.
@@
 Adding object labels leads to consistent performance improvements across all benchmarks, demonstrating their importance in guiding the vision-language model toward accurate grounding.
-Appendix C Additional Quantitative Results
+Appendix C Ablation Study on Prompt Design for Object-Level Descriptions
+To assess how different prompt formulations influence the quality of generated object-level relational descriptions and downstream 3D scene understanding, we compare two designs: a default prompt (Prompt A) that emphasizes relational conciseness, and a spatially grounded prompt (Prompt B) that encourages explicit spatial terms and appearance details.
+fig.˜4 presents qualitative examples generated using Prompt A (Default).
+This prompt directs the vision-language model (LLaVA-1.5) to describe relational context in a concise, human-like way, without placing heavy emphasis on precise spatial markers.
+As shown, descriptions tend to mention object co-occurrence and proximity in natural, readable sentences.
+For example, “The curtain is covering the window, and it is also close to a table, a cabinet, and a desk.” While this phrasing lacks precise positional anchoring, it aligns with how humans intuitively describe contextual relevance.
+In contrast, fig.˜5 illustrates the results of Prompt B (Spatially Focused).
+This prompt explicitly encourages the use of geometric relations (“on the left,” “in front of,” “behind”) and appearance details (“white,” “rectangular”), resulting in descriptions that are shorter but more spatially grounded.
+For instance, “The desk is located in a corner of the room… the window is above the desk… the cabinet is in front of the desk,” offers a clearer positional context but a less nuanced interpretation of function or co-usage.
+To better understand how different prompts influence downstream model performance, table˜8 provides the full text of each design, and table˜9 summarizes quantitative results across five 3D vision-language tasks.
+While Prompt B encourages explicit spatial expressions (e.g., “to the left of,” “in front of”) and produces shorter sentences, it tends to focus narrowly on positional details, omitting functional or contextual cues.
+This results in descriptions that are more rigid but less informative overall.
+In contrast, Prompt A (Ours) generates a richer relational context with broader object co-occurrence and usage clues.
+As shown, Prompt A consistently outperforms Prompt B, suggesting that general, semantically rich descriptions better support multimodal reasoning than strictly spatial ones.
+Prompt Version Prompt Text
+Prompt A (Default) Describe clearly and briefly the relationships between the <key_object> in the scene and nearby objects (<other_obj1>, <other_obj2>, ...).
+Do not describe objects you cannot see.
+Do not describe green labels.
+Prompt B (Spatially Focused) Based on the image, describe both the appearance and spatial relationships of the <key_object> in relation to nearby visible objects (<other_obj1>, <other_obj2>, ...).
+Include visual details like color, shape, size, or texture of the <key_object>, and explain precisely how it is positioned relative to nearby visible objects (<other_obj1>, <other_obj2>, ...) using terms such as ’on the left’, ’next to’, ’under’, ’in front of’, ’behind’, or ’on top of’.
+Only refer to what is clearly visible.
+Do not mention green text labels or objects not shown in the image."
+Table 8: Comparison of prompt designs used for generating object-level relational descriptions with LLaVA-1.5.
+Prompt A is our default, concise formulation emphasizing relational grounding.
+Prompt B explicitly encourages spatial terms (e.g., “left,” “in front of”) and detailed appearance cues.
+Prompt Design ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
+Acc@0.5 F1@0.5 C@0.5 CIDEr EM
+Prompt B 51.4 55.1 74.1 92.3 55.2
+Prompt A (Ours) 51.8 55.1 77.2 93.7 55.7
+Table 9: Downstream performance using different prompts for generating object-level descriptions.
+Prompt B emphasizes spatial precision, while Prompt A (ours) encourages concise, general relational reasoning.
+Despite lacking explicit directional terms, Prompt A outperforms or matches Prompt B, suggesting that overly specific spatial descriptions may omit broader contextual signals useful for multimodal understanding.
+Appendix D Ablation Study on Description Generator Choice
+We further analyze the effect of the model used to generate object-level descriptions.
+While our main experiments adopt Vicuna-7B as the trainable backbone (to ensure fair comparison with prior methods such as Chat-Scene and 3DGraphLLM), the relational text can, in principle, be generated by any frozen captioner.
+table˜10 compares two options: generating descriptions with Vicuna-7B versus with LLaVA-1.5.
+Results show that replacing Vicuna with LLaVA-1.5 as the description generator improves downstream performance, especially on language-intensive tasks (e.g., +2.0 CIDEr on Scan2Cap, +2.1 CIDEr on ScanQA).
+This suggests that our framework is flexible with respect to the choice of description generator, and benefits from relation-dense captions produced by stronger multimodal models.
+VLM ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
+Acc@0.5 F1@0.5 C@0.5 CIDEr EM
+Vicuna-7B 51.8 55.0 75.2 91.6 55.6
+LLaVa-1.5 (Ours) 51.8 55.1 77.2 93.7 55.7
+Table 10: Impact of the choice of description generator.
+We compare object-level relational text produced by Vicuna-7B and by LLaVA-1.5 (used as frozen captioners).
+Using LLaVA-1.5 yields stronger downstream results.
+Appendix E Additional Qualitative Results and Failure Cases
+Additional Qualitative Examples
+To further demonstrate the strengths of Descrip3D, we present additional qualitative comparisons of both question answering and object grounding tasks in fig.˜6.
+In the QA task fig.˜6(a), Descrip3D produces accurate answers in cases where Chat-Scene fails due to limited spatial awareness or insufficient contextual cues.
+For example, in the first question, while Chat-Scene incorrectly places the "single seat sofa" behind the brown chair, Descrip3D correctly identifies it as “in the corner of the room,” grounded by relational language.
+Similarly, Descrip3D succeeds in localizing queried objects such as the laptop and chairs based on complex object-to-object references, demonstrating its enhanced relational understanding.
+In the grounding task fig.˜6(b), Descrip3D resolves ambiguous references more reliably.
+For example, given a query like “the black couch next to a tall shelf and a fan,” Descrip3D identifies the correct object using spatial and contextual signals provided by the object descriptions.
+These results emphasize Descrip3D’s ability to perform robust reasoning in cluttered indoor environments where visual and geometric cues alone may be insufficient.
+Failure Case Analysis
+Despite its improved performance, Descrip3D is not immune to errors.
+fig.˜7 illustrates several representative failure cases in both QA (fig.˜7(a)) and grounding (fig.˜7(b)).
+In question answering, a common failure mode arises in counting tasks where performance is limited by upstream 3D detection accuracy.
+For instance, when the detector undercounts chairs around a table, Descrip3D cannot recover the correct answer solely through textual reasoning.
+Additionally, discrepancies between query phrasing (e.g., “square table”) and detector-generated object names (e.g., “coffee table”) introduce challenges in aligning language inputs with the available descriptions.
+In grounding, failures often occur when key attributes mentioned in the query (e.g., color or material) are missing or omitted in the generated object descriptions.
+As shown in  fig.˜7(b), the system fails to ground “the black cotton pillow” because the corresponding object description lacks explicit mention of its color, resulting in ambiguity during matching.
+These cases highlight limitations in both object detection accuracy and object description completeness, pointing to future directions for improving the robustness and coverage of relational grounding in 3D scene understanding.
+(a) Additional qualitative comparison of question answering.
+Descrip3D correctly answers more challenging questions by leveraging precise spatial relations and context-aware relational cues from its textual descriptions.
+(b) Additional qualitative comparison of grounding.
+Descrip3D provides more accurate object localization by resolving ambiguous object references using context-enhanced descriptions, such as "left of the fan" or “next to another door.”
+Figure 6: Additional Qualitative comparison of 3D scene understanding tasks.
+Descrip3D outperforms Chat-Scene, especially in cases involving complex spatial grounding or multi-object reasoning, due to its use of a dual-level integrated relational textual descriptions that enhance contextual understanding.
+(a) Additional qualitative comparison of question answering.
+Descrip3D fails in cases where the 3D detector misses objects or the query uses object names not aligned with detector output, limiting the effectiveness of textual reasoning.
+(b) Additional qualitative comparison of grounding.
+Failures in grounding are often caused by incomplete object descriptions that omit key attributes (e.g., “black cotton pillow”), which are essential for accurate reference resolution.
+Figure 7: Failure cases of 3D scene understanding tasks.
+While Descrip3D improves contextual reasoning, failure can still occur due to missing or ambiguous descriptions (e.g., color not mentioned) or mismatches between detection names and query expressions, especially in counting or spatial referencing.
+Appendix F Additional Quantitative Results
 We evaluate our method using the standard metrics established in the original papers for each 3D scene-language dataset.
@@
 To complement the main results, we report additional evaluation metrics on the same datasets (ScanRefer, Multi3DRefer, and ScanQA) used in the main paper.
-The results, summarized in Table 8 (ScanRefer), Table 9 (Multi3DRefer), and Table 10 (ScanQA), show our method consistently outperforms prior approaches across grounding and question answering tasks.
+The results, summarized in table˜12 (ScanRefer), table˜13 (Multi3DRefer), and table˜14 (ScanQA), show our method consistently outperforms prior approaches across grounding and question answering tasks.
 On ScanRefer, Descrip3D achieves the highest overall accuracy.
@@
 These results confirm the effectiveness of incorporating object-level textual descriptions through dual-level integration for 3D vision-language tasks.
-Appendix D Prompt Template
+Appendix G Prompt Template
 […]
+Figure 8: IoU distribution comparison between Chat-Scene and Descrip3D on (a) ScanRefer and (b) Multi3DRefer.
+The histograms show the per-sample IoU distributions, with dashed vertical lines indicating the 0.25 and 0.5 thresholds commonly used for visual grounding.
+Compared to Chat-Scene, Descrip3D consistently reduces the concentration of low-IoU cases and increases the density of high-IoU predictions, demonstrating more robust performance across the full distribution.
 […]
-User: What is the ID of the object that matches the description "this is a long table. it is surrounded by chairs"?
-Assistant: <OBJ023>
+User: What is the ID of the object that matches the description "this is a brown chair.
+It is to the right of another chair near the end of the table."?
+[Generated description (may be noisy)] <OBJ002>: there is a <OBJ002> positioned next to a <OBJ003>…
+Assistant: <OBJ002>
 Table 11: Prompt template used during training and evaluation.
@@
 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
-ScanRefer [4] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
-TGNN [24] AAAI21 68.61 56.80 29.84 23.18 37.37 29.70
-SAT [51] ICCV21 73.21 50.83 37.64 25.16 44.54 30.14
-InstanceRefer [55] ICCV21 75.72 64.66 29.41 22.99 38.40 31.08
-3DVG-Transformer [59] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
-MVT [25] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
-3D-SPS [32] CVPR22 84.12 66.72 40.32 29.82 48.82 36.98
-ViL3DRel [7] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
-3DJCG [3] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
-D3Net [5] ECCV22 72.04 30.05 37.87
-BUTD-DETR [26] ECCV22 84.2 66.3 46.6 35.1 52.2 39.8
-HAM [6] ArXiv22 79.24 67.86 41.46 34.03 48.79 40.60
-3DRP-Net [43] EMNLP23 83.13 67.74 42.14 31.95 50.10 38.90
-3D-VLP [27] CVPR23 84.23 64.61 43.51 33.41 51.41 39.46
-EDA [47] CVPR23 85.76 68.57 49.13 37.64 54.59 42.26
-M3DRef-CLIP [58] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
-3D-VisTA [62] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
-ConcreteNet [41] ECCV24 86.40 82.05 42.41 38.39 50.61 46.53
-DORa [46] ArXiv24 52.80 44.80
-Chat-Scene [22] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
+ScanRefer [chen2020scanrefer] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
+TGNN [huang2021text] AAAI21 68.61 56.80 29.84 23.18 37.37 29.70
+SAT [yang2021sat] ICCV21 73.21 50.83 37.64 25.16 44.54 30.14
+InstanceRefer [yuan2021instancerefer] ICCV21 75.72 64.66 29.41 22.99 38.40 31.08
+3DVG-Transformer [zhao20213dvg] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
+MVT [huang2022multi] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
+3D-SPS [luo20223d] CVPR22 84.12 66.72 40.32 29.82 48.82 36.98
+ViL3DRel [chen2022language] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
+3DJCG [cai20223djcg] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
+D3Net [chen2022d3net] ECCV22 72.04 30.05 37.87
+BUTD-DETR [jain2022bottom] ECCV22 84.2 66.3 46.6 35.1 52.2 39.8
+HAM [chen2022ham] ArXiv22 79.24 67.86 41.46 34.03 48.79 40.60
+3DRP-Net [wang20233drp] EMNLP23 83.13 67.74 42.14 31.95 50.10 38.90
+3D-VLP [jin2023context] CVPR23 84.23 64.61 43.51 33.41 51.41 39.46
+EDA [wu2023eda] CVPR23 85.76 68.57 49.13 37.64 54.59 42.26
+M3DRef-CLIP [zhang2023multi3drefer] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
+3D-VisTA [zhu20233d] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
+ConcreteNet [unal2023three] ECCV24 86.40 82.05 42.41 38.39 50.61 46.53
+DORa [wu2024dora] ArXiv24 52.80 44.80
+Chat-Scene [huang2024chat] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
 Descrip3D (Ours) 90.79 83.23 49.62 44.72 57.24 51.84
-Table 8: Performance comparison on the validation set of ScanRefer [4].
+Table 12: Performance comparison on the validation set of ScanRefer [chen2020scanrefer].
 Method Venue ZT w/o D ZT w/ D ST w/o D ST w/ D MT ALL
 F1 F1 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5
-3DVG-Trans+ [59] ICCV21 87.1 45.8 16.7 26.5 25.5
-D3Net (Grounding) [5] ECCV22 81.6 32.5 23.3 35.0 32.2
-3DJCG (Grounding) [3] CVPR22 94.1 66.9 16.7 26.2 26.6
-M3DRef-CLIP [58] ICCV23 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
-Chat-Scene [22] NeurIPS24 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
+3DVG-Trans+ [zhao20213dvg] ICCV21 87.1 45.8 16.7 26.5 25.5
+D3Net (Grounding) [chen2022d3net] ECCV22 81.6 32.5 23.3 35.0 32.2
+3DJCG (Grounding) [cai20223djcg] CVPR22 94.1 66.9 16.7 26.2 26.6
+M3DRef-CLIP [zhang2023multi3drefer] ICCV23 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
+Chat-Scene [huang2024chat] NeurIPS24 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
 Descrip3D (Ours) 92.0 70.4 83.1 75.9 51.4 47.4 49.2 45.2 59.4 55.1
-Table 9: Performance comparison on the validation set of Multi3DRefer [58].
+Table 13: Performance comparison on the validation set of Multi3DRefer [zhang2023multi3drefer].
 Method Venue EM@1 B-1 B-2 B-3 B-4 ROUGE-L METEOR CIDEr
-ScanQA [2] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
-3D-VLP [27] CVPR22 21.65 30.53 21.33 16.67 11.15 34.51 13.53 66.97
-3D-LLM [20] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
-LL3DA [8] CVPR24 13.53 37.31 15.88 76.79
-LEO [23] ICML24 11.5 39.3 16.2 80.0
-Scene-LLM [14] WACV25 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
-Chat-Scene [22] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
+ScanQA [azuma2022scanqa] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
+3D-VLP [jin2023context] CVPR22 21.65 30.53 21.33 16.67 11.15 34.51 13.53 66.97
+3D-LLM [hong20233d] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
+LL3DA [chen2024ll3da] CVPR24 13.53 37.31 15.88 76.79
+LEO [huang2023embodied] ICML24 11.5 39.3 16.2 80.0
+Scene-LLM [fu2024scene] WACV25 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
+Chat-Scene [huang2024chat] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
 Descrip3D (Ours) 22.67 44.36 30.51 22.08 15.70 43.01 19.06 93.71
-Table 10: Performance comparison on the validation set of ScanQA [2].
+Table 14: Performance comparison on the validation set of ScanQA [azuma2022scanqa].
+Appendix H IoU Variance and Distribution
+To provide a more fine-grained view of grounding robustness beyond thresholded metrics (Acc@0.25, Acc@0.50), we report IoU distributions and variance.
+fig.˜8 shows per-sample IoU histograms for ScanRefer and Multi3DRefer, both following the typical bimodal pattern with concentrations near IoU = 0 (failure) and IoU = 1 (successful grounding).
+Compared to Chat-Scene, Descrip3D consistently reduces the density of near-zero IoUs and increases the high-IoU mass.
+The measured variance is stable across datasets: ScanRefer variance $\approx$ 0.19 (std $\approx$ 0.44) and Multi3DRefer variance $\approx$ 0.20 (std $\approx$ 0.44).
+These results demonstrate that improvements are not driven by a handful of outlier cases but reflect consistent robustness across samples.